### v2

# Test-time Adaptive Vision-and-Language Navigation

Abstract
Vision-and-Language Navigation (VLN) has witnessed significant advancements in recent years, largely attributed to meticulously curated datasets and proficiently trained models. Nevertheless, when tested in diverse environments, the trained models inevitably encounter significant shifts in data distribution, highlighting that relying solely on pre-trained and fixed navigation models is insufficient. To enhance models’ generalization ability, test-time adaptation (TTA) demonstrates significant potential in the computer vision field by leveraging unlabeled test samples for model updates. However, simply applying existing TTA methods to the VLN task cannot well handle the adaptability-stability dilemma of VLN models, i.e., frequent updates can result in drastic changes in model parameters, while occasional updates can make the models ill-equipped to handle dynamically changing environments. Therefore, we propose a Fast-Slow Test-Time Adaptation (FSTTA) approach for VLN by performing decomposition-accumulation analysis for both gradients and parameters in a unified framework. Specifically, in the fast update phase, gradients generated during the recent multi-step navigation process are decomposed into components with varying levels of consistency. Then, these components are adaptively accumulated to pinpoint a concordant direction for fast model adaptation. In the slow update phase, historically recorded parameters are gathered, and a similar decomposition-accumulation analysis is conducted to revert the model to a stable state. Extensive experiments show that our method obtains impressive performance gains on four popular benchmarks.
1 Introduction
Developing intelligent agents capable of adhering to human directives remains a significant challenge in embodied AI. Recently, Vision-and-Language Navigation (VLN) [3, 52, 11, 38, 54, 79], which requires an agent to comprehend natural language instructions and subsequently execute proper actions to navigate to the target location, serves as a useful platform for examining the instruction-following ability. Despite tremendous progress has been achieved such as transformer-based sequence-to-sequence learning [25, 9, 11], large-scale training data collection [10, 77], and various reinforcement and imitation learning strategies [74, 17], the navigational capabilities of agents within varied testing environments still warrant further improvement.
Figure 1: (a) Illustration of the data distribution shift between training and testing samples in the VLN Task. (b) Comparison between various TTA strategies on REVERIE [52] validation unseen set using OSR and SR metrics. Among them, ‘DUET’ [11] is the base model, ‘Frequent Update’ means updating at certain intervals within each sample, ‘Stable Update’ refers to initializing with the original base model for each sample and using its best in-sample update interval INT=1. All these strategies adopt TENT [68] for model updates. The results show that overly fast or overly slow TTA fail to achieve significant improvements.
In the VLN task, agents are required to sequentially execute actions contingent upon the evolving environmental cues. Regrettably, owing to disparities in environmental factors, such as distinct room types and objects as shown in Figure 1(a), the trained agents inevitably confront significant shifts in data distribution when applied in practical scenarios [20, 22]. In light of this issue, depending solely on a pre-trained and fixed VLN model is inadequate.
Recently, Test-Time Adaptation (TTA) [39, 40, 51, 71] has been recognized as an effective technique for leveraging unlabeled test samples to update models and address shifts in data distribution. It has garnered notable success across various computer vision tasks, such as image classification [68, 51], segmentation [71, 67], and video classification [42, 80]. For instance, TENT [68] utilizes an entropy minimization objective to update model parameters, thereby enhancing the generalization ability to recognizing test data. Nonetheless, the application of TTA in the realm of VLN remains relatively uncharted. Although prevailing TTA methodologies can be integrated into VLN models with certain alterations, this direct application cannot well handle the adaptability-stability dilemma of models due to the multi-step action-execution nature of VLN. Specifically, in contrast to traditional classification tasks, where a single TTA operation suffices for a test sample, VLN mandates an agent to perform sequential actions within a single test sample. On one hand, while conducting TTA at every (or a few) action steps enables rapid agent adaptation to dynamic environments, frequent model updates may introduce significant model alterations, potentially causing cumulative errors and catastrophic forgetting [71, 50, 60], thus compromising model stability during testing. On the other hand, initializing the same model for stable TTA in each test sample may hinder the model’s ability to adaptively learn experience from historical test samples, thereby impeding its potential for achieving superior performance.  Figure 1(b) shows that both overly fast or overly slow model updates fail to achieve significant performance improvements.
To tackle the above issues, we proposes a Fast-Slow Test-Time Adaptation (FSTTA) method for the VLN tasks. Built upon a unified gradient-parameter decomposition-accumulation framework, our approach consists of a fast update phase and a slow update phase, pursuing a balance between adaptability and stability in model updates. Specifically, with a test-time training objective, such as entropy minimization, we can derive gradients at each action step in the fast update phase. However, due to the unsupervised nature of TTA, these gradients inevitably contain noise information. Using these gradients for model update can interfere with the adaptability, especially when the update is frequently invoked. Therefore, we attempt to find a reliable optimization direction by periodically analyzing the gradients generated during the recent multi-step navigation process. We first establish a local coordinate system to decompose these gradients into components with varying levels of consistency. Subsequently, these components are adaptively accumulated to pinpoint a concordant direction for updating the model. Besides, a gradient variance regularization is incorporated to dynamically adjust the learning rate.
After a certain number of fast updates, the model parameters (also called model state) are recorded. To further mitigate the issues of cumulative errors and catastrophic forgetting that may result from excessively frequent model updates, during the slow update phase, we revert the model to its historical state and conduct a decomposition-accumulation analysis on the parameter variation trajectory for a direct model update. This process is akin to the fast phase but shifts its focus from gradients to the parameters. Both phases are performed alternately during testing to balance the adaptability and stability of the model. As shown in Figure 1(b), the proposed method achieves significant improvement against other model update strategies.
Our contributions can be summarized as follows:
  • $\bullet$
    We investigate the test-time adaptation within the realm of VLN. Our proposed method verifies TTA as a promising and viable avenue for enhancing VLN performance.
  • $\bullet$
    Based on a unified decomposition-accumulation framework for both gradients and parameters, our method ensures swift model adaptability to environmental changes in the short-term fast update phase, while preserves stability throughout the long-term slow update phase.
  • $\bullet$
    Our FSTTA elevates the performance of several leading VLN models across four popular benchmarks. When applied to the notable DUET model [11], our method yields a performance boost of over $5\%$ on the representative discrete/continuous datasets REVERIE/R2R-CE. Furthermore, our method shows superior results compared to other premier TTA techniques.
2 Related Work
Vision-and-Language Navigation (VLN). Recently, VLN tasks have received significant attention in the embodied AI domain and a variety of effective approaches have been proposed [3, 86, 20]. Most existing methods facilitate VLN research by developing powerful techniques for model training, including: (i) Designing advanced network architectures. Sequence-to-sequence framework are the most commonly used one for predicting agent actions from a historical observation series. Early VLN models utilize LSTMs with various attention mechanisms [3, 16, 46, 24] while recent ones [23, 25, 9, 41, 11, 27, 1] resort to the more popular transformer-based methods for multi-modal pre-training. Other architectures are also explored such as graph neural networks [86] and parameter-efficient adapter [55]. (ii) Adopting various training paradigms such as reinforcement and imitation learning [49, 63, 74, 17]. Moreover, to estimate the completeness of instruction following and decide when to conduct backtracking, progress monitoring [45, 85] and back-tracking [30, 46] are also employed to promote training process. (iii) Performing data augmentation for training a stronger model. In recent years, more and more large-scale benchmarks are established via collecting human annotations [33, 86, 57] or creating new environments [52, 10]. Other approaches explore techniques such as mixup and synthesis [43, 29], style transfer [37], or future-view image semantics [36] for data augmentation. (iv) Leveraging additional information for boosting model capacity. Since the goal of VLN is to navigate in photo-realistic environments, there are many kinds of information in the world that can be used such as knowledge [38], 3D scene geometry [44, 78], and landmarks [72, 12]. Though the above approaches have attempted to employ fruitful strategies for training an effective VLN model, they still struggle to adequately address the domain discrepancy between training and testing data. Some studies [55, 54] have tried to use dynamic networks or auxiliary models to cope with varying test data, however, these approaches neither directly minimize the domain gap between the model and test data nor offer dynamic model update strategies.
Test-time Adaptation (TTA). TTA, which allows models to adapt the test data in an online and unsupervised manner, has gained a lot of interest with a host of various approaches proposed in the literature [39, 40, 61, 73, 34]. Existing TTA methods generally rely on batch normalization calibration [48, 84, 19], entropy minimization [68, 51, 50, 64], auxiliary self-supervised task or data regularization [62, 28, 5, 83, 66] to acquire useful information for reducing the domain gap between training and testing data. To stabilize adaptation in continuously changing data distribution, recently, continual test-time adaptation [71, 50, 60, 6, 13, 82], as a more practical setting, has been tentatively explored for addressing the cumulative errors and catastrophic forgetting issues. Until now, test-time adaptation has been preliminarily explored in some sequential data analysis fields such as action recognition [42] and video classification [80]. However, TTA on VLN tasks is yet to be explored.
Gradient-based Methods. Gradients are typically central to modern SGD-based deep learning algorithms. To date, gradient analysis research has predominantly focused on domain generalization (DG) [47, 35, 70, 56, 76, 65], due to the negative impact of conflicting gradients from multiple domains on model optimization. Pioneering works [15, 81, 47] perform gradient surgery at the backpropagation phase via various strategies such as normal plane projection [81] and consensus learning [47]. Other approaches resort to gradient agreement regularization for refining the optimization direction by leveraging sharpness [70] or similarity [58, 56] measurements. Different from the above models that only consider a single-phase gradient surgery in DG, we jointly analyze the gradient-parameter states for a two-phase (fast-slow) TTA in the VLN task.
3 Our Approach
Figure 2: Overall framework of the proposed Fast-Slow Test-Time Adaptation (FSTTA) for VLN tasks. In the fast update phase, taking ‘Sample i’ as an example, the model periodically analyzes the gradients ($\{\bm{g}\}$) generated during the recent multi-step navigation and performs a gradient decomposition-accumulation analysis to pinpoint a concordant direction for model update. After a certain number of fast updates, historical model parameters ($\{\bm{\Theta}\}$) are recorded. In the slow update phase, we revert the model to its historical state and conduct a parameter decomposition-accumulation analysis to learn an optimization path for direct parameter modulation. Note that ‘F’, ‘S’ in the robots means the model parameters after fast and slow updates. ‘$\text{F}_{1}$’ indicates the first fast update within a test sample.
3.1 Preliminaries and Framework Overview
Problem Setup and VLN Base Model. Given a natural language instruction $\bm{I}$, the VLN task requires an agent to find the target viewpoint through the environment by executing a series of actions. During the navigation process, an undirected exploration graph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is progressively constructed, where $\mathcal{V}_{t}$ denotes navigable nodes, $\mathcal{E}_{t}$ indicates the connectivity edges, $t$ is the current timestep. At this moment, the agent receives a panoramic view that contains 36 single images. The panorama is represented by the image features $\mathcal{R}_{t}$ and their object features $\mathcal{O}_{t}$, where these features can be extracted by pre-trained vision transformers (ViT) [14, 11, 38]. To accomplish the instruction, the agent needs to predict probabilities for the currently navigable nodes and select the most possible one as the next movement action. The probabilities can be predicted as:
$\bm{s}_{t}=\phi\left(\bm{I},\mathcal{R}_{t},\mathcal{O}_{t},\mathcal{H}_{t};\bm{\Theta}\right),~{}~{}\bm{s}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|}$ (1)
where $\mathcal{H}_{t}$ indicates the history information that encodes the observed visual features and performed actions [11, 55]. $\phi(\cdot)$ is the VLN base model such as the dual-scale graph transformer [11, 10], $\bm{\Theta}$ is the learnable model parameters.
Framework Overview. In this paper, we devote to adjusting the VLN base model during testing process within an unsupervised manner. Our FSTTA framework is illustrated in  Figure 2. For each sample, at timestep $t$, we employ the commonly adopted entropy minimization objective [68, 51] for test-time adaption, which aims to reduce the entropy of the probabilities over the current navigable nodes:
$\mathcal{L}({\bm{s}}_{t};\bm{\Theta})=-\sum\nolimits_{i}{\bm{s}}_{t,i}\log({\bm{s}}_{t,i}).$ (2)
During the optimization of the above objective, gradients are back-propagated for updating the model’s parameters. However, updating the whole base model is computationally infeasible. As a result, we only consider a small portion of the model parameters for gradient calculation. Since affine parameters in normalization layers capturing data distribution information, numerous TTA methods opt to update these parameters for adaption [68, 51, 39]. In this paper, we employ the model’s final a few layer-norm operations for TTA and maintain other parameters frozen. For brevity, we still use the symbol $\bm{\Theta}$ to represent these parameters to be updated, $\bm{\Theta}\in\mathbb{R}^{D}$. Targeting at fully leveraging the gradient and parameter information, under a unified decomposition-accumulation analysis framework, we propose an effective two-phase adaptation for fast and slow model updates.
3.2 Fast Update via Gradient Analysis
At timestep $t$ in the navigation process, the agent is required to select an action (navigable node) by using the predicted score ${\bm{s}}_{t}$. With this score, we can calculate the TTA loss (Eq. (2)) and then derive the gradient of the model parameters $\bm{\Theta}$ as: $\bm{g}_{t}=\nabla\mathcal{L}({\bm{s}}_{t};\bm{\Theta})$, $\bm{g}_{t}\in\mathbb{R}^{D}$. Traditional TTA methods conduct adaptation independently at each time step, which can exacerbate the issue of cumulative errors [50, 60], particularly in the VLN process that requires frequent action execution. Therefore, we propose to conduct a gradient decomposition-accumulation analysis, wherein we periodically analyze the gradients generated during the recent multi-step navigation process and identify a concordant direction for an iteration of model update.
Gradient Decomposition-Accumulation. During navigation, as shown in Figure 2, we perform model update every $M$ action steps. For the $j$-th update, gradients from previous $M$ steps are collected as $\bm{G}_{j}=\{\widetilde{\bm{g}}_{j,m}\}_{m=1}^{M}$, where $\bm{G}_{j}\in\mathbb{R}^{M\times D}$, $\widetilde{\bm{g}}_{j,m}$ indicates the $t$-th gradient $\bm{g}_{t}$ when $t=M(j-1)+m$. Note that these gradients determines the learning direction of our VLN model, and a simple strategy to compute this direction is to take their average $\bar{\bm{g}}_{j}=1/M\sum_{m}\widetilde{\bm{g}}_{j,m}$; however, this inevitably introduce step-specific noise. To avoid the issue, we aim to find a concordant direction among these gradients. We first establish a local coordinate system with $D$ orthogonal and unit axes (bases) $\bm{U}_{j}=\{{\bm{u}_{j,d}}^{\mathsf{T}}\}_{d=1}^{D}\in\mathbb{R}^{D\times D}$ for gradient decomposition, where each gradient can be approximatively linearly represented by these bases. Intuitively, the axes along which the gradients exhibit the higher variance after projection represent the directions of gradients with the lower consistency. These directions have the potential to introduce interference in determining a model update direction. Therefore, it is advisable to reduce the projection of gradients in these directions. To solve the bases $\bm{U}_{j}$, we can utilize singular value decomposition (SVD) as follows:
$\lambda_{j,d},\bm{u}_{j,d}=\textbf{SVD}_{d}\left(\frac{1}{M-1}{\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right),$ (3)
where $\hat{\bm{G}}_{j}$ is the centered gradient matrix by removing the mean from $\bm{G}_{j}$. The $m$-th row vector in $\hat{\bm{G}}_{j}$ reflects the deviation between $\widetilde{\bm{g}}_{j,m}$ and the average gradient $\bar{\bm{g}}_{j}$. $\lambda_{j,d}$, $\bm{u}_{j,d}$ denote the $d$-th largest eigenvalue and the corresponding eigenvector. Motivated by the principle component analysis [59], it is obvious that a larger $\lambda_{j,d}$ corresponds to a higher variance of the gradient projection length $\bm{G}_{j}\bm{u}_{j,d}$ and vice versa. Hence, we can derive a concordant gradient by adaptively aggregating the gradients’ components on all the axes by considering different eigenvalue (importance):
$\nabla_{j}^{(fast)}=\sum\nolimits_{d=1}^{D}\Phi_{d}(\lambda_{j,d})\cdot<\bar{\bm{g}}_{j},\bm{u}_{j,d}>\bm{u}_{j,d},$ (4)
where the last term denotes the projected component of the averaged gradient $\bar{\bm{g}}_{j}$ on to the $d$-th axis. $\Phi_{d}(\cdot)$ is referred as the adaptive coefficient for accumulating all the components, which is simply defined as $\Phi_{d}(\lambda_{j,d})=1/\lambda_{j,d}$, reflecting the importance of various axes. Notably, when removing the coefficient, $\nabla_{j}^{(fast)}$ is degenerated into $\bar{\bm{g}}_{j}$, which is used in regular gradient descent approaches.
Based on Eq. (4), a concordant optimization direction is established by enhancing the components that are convergent among $\{\widetilde{\bm{g}}_{j,m}\}_{m=1}^{M}$ and suppressing those divergent ones. However, the introduction of $\Phi_{d}(\cdot)$ makes the length of $\nabla_{j}^{(fast)}$ uncontrollable. Therefore, we calibrate its length to $\left\|\bar{\bm{g}}_{j}\right\|_{2}$, which encodes the gradient length from the last three time steps, for a more reasonable model update:
$\nabla_{j}^{(fast)}\leftarrow(\nabla_{j}^{(fast)}\|\bar{\bm{g}}_{j}\|_{2})/\|\nabla_{j}^{(fast)}\|_{2}$ (5)
With $\nabla^{(fast)}$, we can perform fast model update by setting a learning rate $\gamma^{(fast)}$. Although traditional methods employ a fixed learning rate during optimization, such a setting might hinder model convergence, i.e., small learning rates slow down convergence while aggressive learning rates prohibit convergence [4]. Since fast updates are frequently invoked during navigation, relying on a fixed learning rate is sub-optimal. Therefore, we propose to dynamically adjust learning rate throughout the fast update phase.
Dynamic Learning Rate Scaling. Different from varying the learning rate through optimizer or scheduler, we argue for a scaling method that leverages gradient agreement information in historical steps to dynamically adjust the speed of model update. Current gradient alignment strategies typically impose direct constraints on the gradients [58, 56], which are not suitable for our framework as they undermine the gradient decomposition-accumulation process. Given that the second-order information (variance) has been demonstrated to be more effective than the first-order information (mean) in gradient agreement learning [56], we directly utilize the trace of the gradient covariance matrix, $\text{Tr}\left(1/(M-1){\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right)$, for scaling. Note that the trace is equal to the sum of eigenvalues $\sigma_{j}=\sum_{d}\lambda_{j,d}$. Here, when $\sigma_{j}$ deviates significantly from the historical variance, we assign a smaller learning rate, and vice versa:
$\gamma^{(fast)}_{j}=\text{Trunc}\left(1+\tau-|\sigma_{j}-\bar{\sigma}|\right)\cdot\hat{\gamma}^{(fast)},$ (6)
where $\text{Trunc}(\cdot)$ is the truncation function that truncates the input to the interval $[a,b]$. $\tau$ is a threshold and $\hat{\gamma}^{(fast)}$ is the base learning rate. The historical variance $\bar{\sigma}$ is updated as $\bar{\sigma}\leftarrow\rho\bar{\sigma}+(1-\rho)\sigma_{j}$ and maintained for all samples throughout the test stage, $\rho$ is the update momentum.
Model Update. With the above gradient and learning rate, we can perform the $j$-th fast model update:
$\bm{\Theta}_{j}=\bm{\Theta}_{j-1}-\gamma_{j}^{(fast)}\cdot\nabla_{j}^{(fast)},$ (7)
where the subscript of $\bm{\Theta}$ indicates the index of model update in the current test sample.
3.3 Slow Update via Parameter Analysis
In the fast update phase, although we obtain concordant optimization directions, the frequent parameter updates may still dramatically change the VLN model. To maintain the stability of the VLN model during long-term usage, we revert the model to its historical states recorded in the fast update phase, and conduct a decomposition-accumulation analysis on the parameter variation trajectory for direct parameter modulation. The slow update phase shares the core formulation with the fast phase, but shifts the focus from gradients to the model parameters themselves.
Parameter Decomposition-Accumulation. Following the completion of the fast update phase on the $o$-th test sample, the model state (parameters) is recorded as $\bm{\Theta}_{o,J_{o}}$, where $J_{o}$ denotes the final fast update step on this sample, and the subscript $o$ has been omitted in the previous section. We then treat these historical states as a parameter variation trajectory to facilitate stable model updates. As shown in the right part of Figure 2, the slow model update is invoked every $N$ samples. For the $l$-th update, historical model states are collected as $\bm{M}_{l}=\{\widetilde{\bm{\Theta}}_{l,n}\}_{n=0}^{N}$, where $\bm{M}_{l}\in\mathbb{R}^{(N+1)\times D}$, $\widetilde{\bm{\Theta}}_{l,n}$ indicates the $o$-th model state $\bm{\Theta}_{o,J_{o}}$ when $o=N(l-1)+n$ and $n\neq 0$. $\widetilde{\bm{\Theta}}_{l,0}$ indicates the model state produced by the previous slow update, and we use it interchangeably with $\bm{\Theta}^{(l-1)}$ in the following. Note that in the slow update phase, we additionally incorporate $\bm{\Theta}^{(l-1)}$ from the previous update for analysis since it serves as a starting reference point for direct parameter modulation.
Table 1: Comparison on REVERIE. Results better than the base model are highlighted in bold.
Table 2: Results on R2R dataset.
Methods REVERIE Val Unseen REVERIE Test Unseen
TL $\downarrow$ OSR SR SPL RGS RGSPL TL $\downarrow$ OSR SR SPL RGS RGSPL
Human - - - - - - 21.18 86.83 81.51 53.66 77.84 51.44
Seq2Seq [3] [CVPR18] 11.07 8.07 4.20 2.84 2.16 1.63 10.89 6.88 3.99 3.09 2.00 1.58
RCM [74] [CVPR19] 11.98 14.23 9.29 6.97 4.89 3.89 10.60 11.68 7.84 6.67 3.67 3.14
SMNA [45] [ICLR19] 9.07 11.28 8.15 6.44 4.54 3.61 9.23 8.39 5.80 4.53 3.10 2.39
FAST [52] [CVPR20] 45.28 28.20 14.40 7.19 7.84 4.67 39.05 30.63 19.88 11.61 11.28 6.08
Airbert [21] [ICCV21] 18.71 34.51 27.89 21.88 18.23 14.18 17.91 34.20 30.28 23.61 16.83 13.28
HAMT [9] [NeurIPS21] 14.08 36.84 32.95 30.20 18.92 17.28 13.62 33.41 30.40 26.67 14.88 13.08
HOP [53] [CVPR22] 16.46 36.24 31.78 26.11 18.85 15.73 16.38 33.06 30.17 24.34 17.69 14.34
LANA [75] [CVPR23] 23.18 52.97 48.31 33.86 32.86 22.77 18.83 57.20 51.72 36.45 32.95 22.85
BEVBert [1] [ICCV23] - 56.40 51.78 36.37 34.71 24.44 - 57.26 52.81 36.41 32.06 22.09
BSG [44] [ICCV23] 24.71 58.05 52.12 35.59 35.36 24.24 22.90 62.83 56.45 38.70 33.15 22.34
GridMM [78] [ICCV23] 23.20 57.48 51.37 36.47 34.57 24.56 19.97 59.55 55.13 36.60 34.87 23.45
DUET [11] [CVPR22] 22.11 51.07 46.98 33.73 32.15 23.03 21.30 56.91 52.51 36.06 31.88 22.06
DUET-FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 21.52 58.44 53.40 36.43 32.99 22.40
HM3D [10] [ECCV22] 22.13 62.11 55.89 40.85 36.58 26.76 20.87 59.81 53.13 38.24 32.69 22.68
HM3D-FSTTA 22.37 63.74 57.02 41.41 36.97 26.55 21.90 63.68 56.44 39.58 34.05 23.04
Methods R2R Val Unseen
TL $\downarrow$ NE $\downarrow$ SR SPL
Seq2Seq [3] 8.39 7.81 22 -
RCM [74] 11.46 6.09 43 -
SMNA [45] - 5.52 45 32
EnvDrop [63] 10.70 5.22 52 48
AirBert [21] 11.78 4.10 62 56
HAMT [9] 11.46 3.65 66 61
GBE [86] - 5.20 54 43
SEvol [7] 12.26 3.99 62 57
HOP [53] 12.27 3.80 64 57
BEVBert [1] 14.55 2.81 75 64
LANA [75] 12.00 - 68 62
BSG [44] 14.90 2.89 74 62
GridMM [78] 13.27 2.83 75 64
DUET [11] 13.94 3.31 72 60
DUET-FSTTA 14.64 3.03 75 62
HM3D [10] 14.29 2.83 74 62
HM3D-FSTTA 14.86 2.71 75 63
Table 2: Results on R2R dataset.
Similar to the fast update phase, the centered parameter matrix $\hat{\bm{M}}_{l}$ can be constructed, where the $n$-th row vector in it reflects the deviation between $\widetilde{\bm{\Theta}}_{l,n}$ and the averaged historical parameter $\bar{\bm{\Theta}}_{l}=1/(N+1)\sum_{n}\widetilde{\bm{\Theta}}_{l,n}$. With $\hat{\bm{M}}_{l}$, we can obtain the following eigenvalues and eigenvectors: $\epsilon_{l,d},\bm{z}_{l,d}=\textbf{SVD}_{d}(1/N\cdot{\hat{\bm{M}}}_{l}^{\mathsf{T}}\hat{\bm{M}}_{l})$, where a larger $\epsilon_{l,d}$ corresponds to a higher variance of the parameter projection length ${\bm{M}}_{l}\bm{z}_{l,d}$ and vice versa. $\bm{Z}_{l}=\{{\bm{z}_{l,d}}^{\mathsf{T}}\}_{d=1}^{D}$ depicts the local coordinate system where each axis depicts the direction of parameter variation. Intuitively, the principal axes (with larger eigenvalues) delineate the primary directions of historical parameter variation, while minor axes (with smaller eigenvalues) often encompass noise [76]. To find a more reliable optimization path to traverse the trajectory of primary parameter changes, we pay more attention on the axis with the larger variance. Since there is no silver bullet to learning an optimization direction with only parameters, a reference direction can significantly aid in guiding the model towards a local optimal. Here, we leverage the parameter variations to calculate the reference direction:
$\bm{h}_{l}=\frac{1}{\sum_{i=0}^{N-1}q^{i}}\sum\nolimits_{n=1}^{N}q^{N-n}\cdot(\widetilde{\bm{\Theta}}_{l,0}-\widetilde{\bm{\Theta}}_{l,n}),$ (8)
where the hyper-parameter $q\in(0,1)$, which assigns larger weight to the more recent parameter deviations as they encapsulate richer sample information. Then, we calculate the optimization path (gradient) in the slow update phase as:
$\nabla_{l}^{(slow)}=\sum\nolimits_{d}\Psi_{d}(\bm{\epsilon}_{l},\bm{h}_{l})\cdot\text{sign}\left(<\bm{h}_{l},\bm{z}_{l,d}>\right)\bm{z}_{l,d},$ (9)
where the use of sign function $\text{sign}(\cdot)$ is to force the axes to be positively related to the reference direction $\bm{h}_{l}$. Notably, different from Eq. (4) that uses the projected components on each axis for estimating an optimization direction, here we only utilize the axes themselves for deriving $\nabla_{l}^{(slow)}$. The reason is that these axes depict the parameter variation direction, which can be directly used for estimating gradients. $\Psi_{d}(\cdot)$ is referred as the adaptive coefficient for accumulating all the axes (optimization directions), defined as:
$\Psi_{d}(\bm{\epsilon}_{l},\bm{h}_{l})=\frac{\epsilon_{l,d}\cdot\|\bm{h}_{l}\|_{2}}{\|\bm{\epsilon}_{l}\|_{2}},$ (10)
where the L2-normalization is performed on eigenvalues to convey different relative importance of axes. Besides, the norm of the reference direction is utilized to automatically tuning the magnitude of the analyzed gradient. In contrast to $\Phi_{d}(\cdot)$ in gradient analysis, $\Psi_{d}(\cdot)$ highlights those axes with high variation due to the different characteristics of gradients and parameters in model optimization.
Model Update. With $\nabla_{l}^{(slow)}$, we can perform the $l$-th slow model update as follows:
$\bm{\Theta}^{(l)}=\bm{\Theta}^{(l-1)}-\gamma^{(slow)}\cdot\nabla_{l}^{(slow)},$ (11)
where $\gamma^{(slow)}$ is learning rate. Since the slow update phase is designed for stable model learning and is not frequently invoked, we employ a fixed learning rate here instead of conducting dynamic learning rate scaling, as done in the fast phase. The updated parameter $\bm{\Theta}^{(l)}$ will be utilized for the subsequent test samples in conjunction with new fast update phases applied to them.
4 Experimental Results
We evaluate FSTTA on four benchmarks: REVERIE [52], R2R [3], SOON [86], and R2R-CE [32] datasets. Experiments and ablation studies show our effectiveness.
4.1 Experimental Setup
Datasets. Four datasets are adopted for conducting our experiments. Among them, REVERIE [52] contains 10,567 panoramic images and 21,702 high-level instructions, focusing on grounding remote target object within 90 buildings. R2R [3] provides step-by-step instructions for navigation in photo-realistic environments, which includes 10,800 panoramic views and 7,189 trajectories. SOON [86] also requires the agent to find the target object with a more detailed description of the goal. It has 3,848 sets of instruction and more than 30K long distance trajectories. R2R-CE [32] is a variant of R2R in continuous environments, where an agent is able to move freely and engage with obstacles. The dataset consists of 16,000 instruction-trajectory pairs, with non-transferrable paths excluded.
Evaluation Metrics. We follow previous approaches [52, 11, 10, 37, 78] and employ the most commonly used metrics for evaluating VLN agents as follows: TL (Trajectory Length), NE (Navigation Error), SR (Success Rate), SPL (Success weighted by Path Length), OSR (Oracle Success Rate), RGS (Remote Grounding Success rate), and RGSPL (RGS weighted by Path Length).
Implementation Details. To better conform to practical scenarios, for all datasets, we set the batch size to 1 during evaluation. Each sample (or each action step) is forward propagated only once during the testing process. We adopt DUET [11] and HM3D [10] as the base models. Since HM3D does not provide training code for R2R-CE dataset, we adopt another state-of-the-art method, BEVBert [1], for TTA. Note that for the base models, in Section 4.2, we report the results obtained from running their official codes. For VLN models equipped with TTA strategies, we run the corresponding experiments 5 times while shuffling the order of the samples and report the average results. In our FSTTA, we only utilize the last four LN layers of base models for model updating, all the feature dimensions of these layers are 768. We set the intervals for fast and slow updates to $M=3$ and $N=4$, the learning rates of the two phases are $\hat{\gamma}^{(fast)}=6\times 10^{-4}$ and $\gamma^{(slow)}=1\times 10^{-3}$. For the dynamic learning rate scaling, we empirically set the threshold $\tau=0.7$ in Eq. (6) and the update momentum $\rho=0.95$ with the truncation interval $[0.9,1.1]$. And the hyper-parameter $q$ in Eq. (8) is set to 0.1. All experiments are conducted on a RTX 3090 GPU.
4.2 Comparison with State-of-the-art VLN Models
REVERIE. Table 2 presents a comparison of our FSTTA against state-of-the-art methods on the REVERIE dataset. Compared with the base models which do not perform test-time adaptation, the proposed method demonstrates favorable performance improvement across most evaluation metrics across the two dataset splits. Specifically, on the validation unseen split, our model exhibits notable advantages over DUET, with improvements of 5.3% on OSR, 7.1% on SR, and 2.7% on SPL. Furthermore, for the recent state-of-the-art method HM3D, our model displays enhanced generalization capabilities on the test unseen split, achieving remarkable improvements over HM3D, including 3.9%, 3.3%, and 1.3% increases on the three metrics. Compared with other state-of-the-arts, our proposed method can achieve superior or comparable performance. These results unequivocally affirm the effectiveness of our fast-slow test time adaptation model, showing the promising potential of TTA in the VLN field. It is noteworthy that none of the prior methods employed a TTA strategy on this task.
R2R. Table 2 shows the comparison results on R2R dataset. Our approach outperforms the base models in most metrics (e.g., $72\%\rightarrow 75\%$ for DUET on SR, $62\%\rightarrow 63\%$ for HM3D on SPL). Notably, from the results of the above two datasets, our method, while enhancing the success rate of VLN, causes a slight increase in the path length (TL). We speculate that a possible reason is that performing TTA online may increase the likelihood of the agent deviating from its original action execution pattern, leading to more exploration or backtracking. This situation is further confirmed in the analysis of various TTA strategies in Table 5.
SOON. The proposed FSTTA establishes new state-of-the-art results across most metrics on this dataset. For instance, as shown in Table 3, on the validation unseen split, our model HM3D-FSTTA achieves SR and SPL of 42.44% and 31.03%, respectively, while the state-of-the-art method GridMM are 37.46% and 24.81%. On the test unseen split, our approach improves the performance of DUET by substantial gains (e.g., $21.42\%\rightarrow 23.23\%$ for SPL).
Table 3: Experimental results on SOON dataset.
Methods Val Unseen Test Unseen
OSR SR SPL RGSPL OSR SR SPL RGSPL
GBE [86] 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
GridMM [78] 53.39 37.46 24.81 3.91 48.02 36.27 21.25 4.15
DUET [11] 50.91 36.28 22.58 3.75 43.00 33.44 21.42 4.17
DUET-FSTTA 52.57 36.53 23.82 3.75 43.44 35.34 23.23 4.52
HM3D [10] 53.22 41.00 30.69 4.06 47.26 40.26 28.09 5.15
HM3D-FSTTA 54.19 42.44 31.03 4.93 48.52 42.02 28.95 5.20
R2R-CE. FSTTA also generalizes well on the continuous environment, i.e., R2R-CE dataset, as shown in Table 4. The results indicate that our approach demonstrates superior or comparable performance against other methods across several metrics.
Table 4: Experimental results on R2R-CE dataset.
Methods Val Unseen Test Unseen
NE $\downarrow$ OSR SR SPL NE $\downarrow$ OSR SR SPL
Seq2Seq [32] 7.37 40 32 30 7.91 36 28 25
CWTP [8] 7.90 38 26 23 - - - -
[…]
Sim2Sim [31] 6.07 52 43 36 6.17 52 44 37
CWP-BERT [26] 5.74 53 44 39 5.89 51 42 36
DREAMW [69] 5.53 49 59 44 5.48 49 57 44
GridMM [78] 5.11 61 49 41 5.64 56 46 39
ETPNav [2] 4.71 65 57 49 5.12 63 55 48
DUET [11] 5.13 55 46 40 5.82 50 42 36
DUET-FSTTA 5.27 58 48 42 5.84 55 46 38
BEVBert [1] 4.57 67 59 50 4.70 67 59 50
BEVBert-FSTTA 4.39 65 60 51 5.45 69 60 50
Table 5: Experimental results for different TTA strategies.
Methods REVERIE Val Unseen Time(ms)
TL $\downarrow$ OSR SR SPL RGS RGSPL
DUET [11] 22.11 51.07 46.98 33.73 32.15 23.03 104.84
 + EATA [50] 23.41 52.09 47.40 33.46 32.09 22.65 133.12
[…]
 + NOTE [19] 23.15 52.85 48.28 33.98 32.77 22.98 137.89
 + SAR [51] 23.47 53.26 48.00 33.92 33.49 23.09 145.53
 + Tent [68] 24.05 49.43 46.87 31.90 30.04 20.15 126.91
 + Tent-INT-2 24.24 51.22 48.46 33.67 32.43 21.30 124.02
 + Tent-INT-3 22.52 52.28 48.60 34.65 32.66 23.12 119.34
 + Tent-INT-4 22.59 51.40 48.91 35.06 32.59 22.99 117.26
 + Tent-Stable 22.05 51.43 47.55 33.99 32.34 23.32 129.22
 + FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 135.61
4.3 Results for Different TTA Strategies
Currently, various TTA methods have been adeptly integrated for the dynamic model updates within diverse computer vision tasks, marking significant progress. Although the exploration of TTA’s application within the VLN field remains relatively untapped, the integration of contemporary advanced TTA methodologies into VLN is feasible. Since efficiency is an important evaluation metric for TTA, we provide the average time taken by each method to execute a single instruction for comparison. Obviously, equipping with TTA inevitably incurs additional time costs. For the compared methods, SAR and TENT are the popular entropy minimization models, whereas NOTE, CoTTA, and EATA are state-of-the-art continual TTA methods. The results in Table 5 demonstrate the capability of our proposed FSTTA to blend model performance with testing efficiency. Specifically, on the validation unseen dataset of REVERIE, our method exhibits a discernible enhancement of 6.2% and 2.5% on the SR and SPL metrics compared to the state-of-the-art SAR method, concurrently manifesting a reduction of 7% in testing time. From the results, directly applying existing TTA methods to the VLN task does not lead to significant performance improvements. Furthermore, we investigate different frequencies of updates based on TENT as well as the stable update approach. ‘INT’ represents the update interval, which means averaging the gradient information over a certain interval and then performing an iteration of model update; these results are consistent with those in Figure 1(b). It can be seen that our method still outperforms these strategies with marginally increased time costs.
Table 6: Ablation study on REVERIE dataset.
Module REVERIE Val Unseen
Fast DLR Slow TL $\downarrow$ OSR SR SPL RGS RGSPL
- - - 22.11 51.07 46.98 33.73 32.15 23.03
Tent - - 22.52 52.28 48.60 34.65 32.66 23.12
- - 22.65 53.50 49.74 34.91 33.70 23.36
- 22.43 54.01 49.82 35.34 34.32 23.29
22.14 56.26 54.15 36.41 34.27 23.56
Table 7: Results on validation seen set of REVERIE.
FSTTA REVERIE Val Seen
Unseen Seen TL $\downarrow$ OSR SR SPL RGS RGSPL
- - 13.86 73.86 71.15 63.94 57.41 51.14
- 15.13 75.59 75.48 65.84 58.62 52.23
- 13.40 73.16 71.78 64.18 57.05 51.18
15.11 75.58 74.12 65.53 59.20 52.18
4.4 Further Remarks
We perform ablation studies and other in-depth analysis of FSTTA on the validation unseen set of REVERIE [52].
Ablation Studies of the Proposed FSTTA. In this work, we propose a FSTTA method for vision-and-language navigation, which consists of both fast and slow model update phases. To validate their effectiveness, we progressively integrate the two phases into the baseline DUET model. In addition, we design a baseline variant, which equips DUET with the vanilla TTA objective (TENT [68]) and simply utilize the averaged gradient in an interval (with the same $M$) for fast model updates. Empirical findings from Table 6 illuminate that the integration of fast and slow phases progressively bolsters the base model by 2.8% and 4.3% on the SR metric. Moreover, the dynamic learning rate scaling module (DLR) also contributes to enhancing the model’s performance. Furthermore, our method surpasses the vallina TTA method by a significant margin, showing the consideration of the fast-slow update mechanism is effective.
Will our method experience catastrophic forgetting? For a VLN agent endowed with the TTA capability, it faces the issue of catastrophic forgetting of historical environments and instructions upon continually executing new instructions in new environments. To assess whether our method harbors this issue, we re-evaluate our methods on REVERIE validation seen data. Compared with the base model, as shown in Table 7, we find that: (1) Directly applying FSTTA with the base model on seen data can noticeably enhance performance. (2) After performing FSTTA on the unseen set, the obtained model, when tested directly on the seen dataset without TTA, achieves performance comparable to the base model, confirming that our method does not suffer from catastrophic forgetting. (3) Applying the updated model from unseen set to the seen set with TTA yields the best results. This indicates that our method is effective in accumulating experience from historical test data.
Generalization Testing in More Practical Environments. In the real-world applications, agents might encounter both previously seen and unseen scenarios. In our preceding experiments, we exclusively test on the validation seen and unseen sets separately. To verify the generalizability, we combine the seen and unseen sets into a unified set. Table 8 shows that FSTTA outperforms other TTA methods in effectively managing a variety of testing scenarios.
Qualitative Analysis. Figure 3 provides a visualization of the agent’s instruction execution process, validating that our proposed FSTTA approach can indeed dynamically enhance the VLN performance of the agent during testing.
Table 8: Results of validation unseen & seen on REVERIE dataset.
Methods REVERIE Val Unseen & Seen
TL $\downarrow$ OSR SR SPL RGS RGSPL
DUET [11] 19.18 61.53 57.49 45.66 41.56 34.38
 + Tent [68] 20.23 57.33 54.86 41.90 38.09 32.46
 + EATA [50] 20.29 62.77 57.31 44.59 41.54 34.16
 + SAR [51] 20.52 63.59 57.80 44.72 42.45 34.88
 + FSTTA 20.48 63.36 60.23 47.96 43.58 35.65
Figure 3: Visual results of DUET [11] and FSTTA on REVERIE validation unseen set. Our method arrives the correct endpoint.
5 Conclusions
This paper explores the feasibility of TTA strategies in VLN. We propose a fast-slow test-time adaptation method, which performs decomposition-accumulation analysis for both gradients and parameters, achieving a balance between adaptability and stability. The encouraging performance is validated in extensive experiments. Several limitations of this paper are noteworthy. Firstly, our approach focuses on adapting normalization layers within the trained model. While normalization layers are widely employed in deep learning, there are still a few methods that do not utilize these settings. One viable approach to address this issue is to introduce additional normalization layers to the corresponding models and retrain them using the training data. In the future, we will also explore how our model can update other types of layers. Secondly, the VLN task itself is a cross-modal learning task. However, our TTA process does not explicitly consider this information. We plan to consider cross-modal TTA in the future. Thirdly, compared to the base model, the introduction of TTA inevitably incurs additional computational cost, which is a direction for future improvement. Finally, the frequencies of fast and slow updates are fixed and periodic. Adaptive update invocation strategies is worthy of consideration.

### v3

@@
-Test-time Adaptive Vision-and-Language Navigation
+Fast-Slow Test-Time Adaptation for Online Vision-and-Language Navigation
 Abstract
-Vision-and-Language Navigation (VLN) has witnessed significant advancements in recent years, largely attributed to meticulously curated datasets and proficiently trained models.
-Nevertheless, when tested in diverse environments, the trained models inevitably encounter significant shifts in data distribution, highlighting that relying solely on pre-trained and fixed navigation models is insufficient.
-To enhance models’ generalization ability, test-time adaptation (TTA) demonstrates significant potential in the computer vision field by leveraging unlabeled test samples for model updates.
-However, simply applying existing TTA methods to the VLN task cannot well handle the adaptability-stability dilemma of VLN models, i.e., frequent updates can result in drastic changes in model parameters, while occasional updates can make the models ill-equipped to handle dynamically changing environments.
-Therefore, we propose a Fast-Slow Test-Time Adaptation (FSTTA) approach for VLN by performing decomposition-accumulation analysis for both gradients and parameters in a unified framework.
-Specifically, in the fast update phase, gradients generated during the recent multi-step navigation process are decomposed into components with varying levels of consistency.
-Then, these components are adaptively accumulated to pinpoint a concordant direction for fast model adaptation.
-In the slow update phase, historically recorded parameters are gathered, and a similar decomposition-accumulation analysis is conducted to revert the model to a stable state.
+The ability to accurately comprehend natural language instructions and navigate to the target location is essential for an embodied agent.
+Such agents are typically required to execute user instructions in an online manner, leading us to explore the use of unlabeled test samples for effective online model adaptation.
+However, for online Vision-and-Language Navigation (VLN), due to the intrinsic nature of inter-sample online instruction execution and intra-sample multi-step action decision, frequent updates can result in drastic changes in model parameters, while occasional updates can make the model ill-equipped to handle dynamically changing environments.
+Therefore, we propose a Fast-Slow Test-Time Adaptation (FSTTA) approach for online VLN by performing joint decomposition-accumulation analysis for both gradients and parameters in a unified framework.
 Extensive experiments show that our method obtains impressive performance gains on four popular benchmarks.
+Code is available at https://github.com/Feliciaxyao/ICML2024-FSTTA.
+Machine Learning, ICML
 1 Introduction
 Developing intelligent agents capable of adhering to human directives remains a significant challenge in embodied AI.
-Recently, Vision-and-Language Navigation (VLN) [3, 52, 11, 38, 54, 79], which requires an agent to comprehend natural language instructions and subsequently execute proper actions to navigate to the target location, serves as a useful platform for examining the instruction-following ability.
-Despite tremendous progress has been achieved such as transformer-based sequence-to-sequence learning [25, 9, 11], large-scale training data collection [10, 77], and various reinforcement and imitation learning strategies [74, 17], the navigational capabilities of agents within varied testing environments still warrant further improvement.
-Figure 1: (a) Illustration of the data distribution shift between training and testing samples in the VLN Task.
-(b) Comparison between various TTA strategies on REVERIE [52] validation unseen set using OSR and SR metrics.
-Among them, ‘DUET’ [11] is the base model, ‘Frequent Update’ means updating at certain intervals within each sample, ‘Stable Update’ refers to initializing with the original base model for each sample and using its best in-sample update interval INT=1.
-All these strategies adopt TENT [68] for model updates.
+Recently, Vision-and-Language Navigation (VLN) (Anderson et al., 2018; Qi et al., 2020; Chen et al., 2022d; Yang et al., 2023), which requires an agent to comprehend natural language instructions and subsequently perform proper actions to navigate to the target location, serves as a useful platform for examining the instruction-following ability.
+Figure 1: (a) Illustration of online VLN.
+(b) Comparison between TTA strategies on REVERIE (Qi et al., 2020) validation unseen set using SPL and SR metrics.
+‘DUET’ (Chen et al., 2022d) is the base model, ‘Frequent Update’ means updating at certain intervals within each sample, ‘Stable Update’ refers to initializing with the original base model for each sample and using its best intra-sample update interval INT=1.
+All these strategies adopt TENT (Wang et al., 2021) for model updates.
 The results show that overly fast or overly slow TTA fail to achieve significant improvements.
-In the VLN task, agents are required to sequentially execute actions contingent upon the evolving environmental cues.
-Regrettably, owing to disparities in environmental factors, such as distinct room types and objects as shown in Figure 1(a), the trained agents inevitably confront significant shifts in data distribution when applied in practical scenarios [20, 22].
-In light of this issue, depending solely on a pre-trained and fixed VLN model is inadequate.
-Recently, Test-Time Adaptation (TTA) [39, 40, 51, 71] has been recognized as an effective technique for leveraging unlabeled test samples to update models and address shifts in data distribution.
-It has garnered notable success across various computer vision tasks, such as image classification [68, 51], segmentation [71, 67], and video classification [42, 80].
-For instance, TENT [68] utilizes an entropy minimization objective to update model parameters, thereby enhancing the generalization ability to recognizing test data.
-Nonetheless, the application of TTA in the realm of VLN remains relatively uncharted.
-Although prevailing TTA methodologies can be integrated into VLN models with certain alterations, this direct application cannot well handle the adaptability-stability dilemma of models due to the multi-step action-execution nature of VLN.
-Specifically, in contrast to traditional classification tasks, where a single TTA operation suffices for a test sample, VLN mandates an agent to perform sequential actions within a single test sample.
-On one hand, while conducting TTA at every (or a few) action steps enables rapid agent adaptation to dynamic environments, frequent model updates may introduce significant model alterations, potentially causing cumulative errors and catastrophic forgetting [71, 50, 60], thus compromising model stability during testing.
+In practical applications, a trained VLN agent is required to execute user instructions in various environments at different times in an online manner, as depicted in Figure 1(a).
+However, owing to disparities in environmental factors, such as distinct room types and objects, the trained agent inevitably confront data discrepancies during online testing (Gu et al., 2022; Guhur et al., 2021).
+This raises an important question: can an agent accumulate experience and enhance its capabilities while executing instructions?
+However, due to the lack of annotation, updating the model via supervised online learning is impractical.
+In addition, other learning paradigms like unsupervised domain adaptation or semi-supervised learning are also infeasible, given considerations for the issues of execution efficiency and privacy protection.
+Recently, online Test-Time Adaptation (TTA) (Liang et al., 2023; Niu et al., 2023; Wang et al., 2022a) has been recognized as an effective technique of online model updating by leveraging unlabeled test samples.
+Although prevailing TTA methods can be integrated into VLN models with certain alterations, this direct application cannot well handle the adaptability-stability dilemma of models due to the intrinsic nature of inter-sample online instruction execution and intra-sample multi-step action decision.
+Specifically, unlike traditional classification tasks where a single TTA operation is sufficient for each test sample, as illustrated in Figure 1(a), online VLN requires an agent to perform a sequence of actions within a single test sample and to execute various samples (instructions) in an online manner.
+On one hand, while conducting TTA at every (or a few) action steps enables rapid agent adaptation to dynamic environments, frequent model updates may introduce significant model alterations, potentially causing cumulative errors and catastrophic forgetting (Wang et al., 2022a; Niu et al., 2022; Song et al., 2023), thus compromising model stability during testing.
 On the other hand, initializing the same model for stable TTA in each test sample may hinder the model’s ability to adaptively learn experience from historical test samples, thereby impeding its potential for achieving superior performance.
 Figure 1(b) shows that both overly fast or overly slow model updates fail to achieve significant performance improvements.
-To tackle the above issues, we proposes a Fast-Slow Test-Time Adaptation (FSTTA) method for the VLN tasks.
-Built upon a unified gradient-parameter decomposition-accumulation framework, our approach consists of a fast update phase and a slow update phase, pursuing a balance between adaptability and stability in model updates.
-Specifically, with a test-time training objective, such as entropy minimization, we can derive gradients at each action step in the fast update phase.
+To tackle the above issues, we propose a Fast-Slow Test-Time Adaptation (FSTTA) method for online VLN tasks.
+Built upon a unified gradient-parameter decomposition-accumulation framework, our approach consists of a fast update phase and a slow update phase, pursuing a balance between adaptability and stability in online model updating.
+Specifically, with a test-time training objective, such as entropy minimization (Wang et al., 2021), we can derive gradients at each action step in the fast update phase.
 However, due to the unsupervised nature of TTA, these gradients inevitably contain noise information.
 Using these gradients for model update can interfere with the adaptability, especially when the update is frequently invoked.
-Therefore, we attempt to find a reliable optimization direction by periodically analyzing the gradients generated during the recent multi-step navigation process.
-We first establish a local coordinate system to decompose these gradients into components with varying levels of consistency.
-Subsequently, these components are adaptively accumulated to pinpoint a concordant direction for updating the model.
-Besides, a gradient variance regularization is incorporated to dynamically adjust the learning rate.
+Therefore, we establish a local coordinate system to find a reliable optimization direction by periodically analyzing the gradients generated during the recent multi-step navigation process.
 After a certain number of fast updates, the model parameters (also called model state) are recorded.
 To further mitigate the issues of cumulative errors and catastrophic forgetting that may result from excessively frequent model updates, during the slow update phase, we revert the model to its historical state and conduct a decomposition-accumulation analysis on the parameter variation trajectory for a direct model update.
-This process is akin to the fast phase but shifts its focus from gradients to the parameters.
 Both phases are performed alternately during testing to balance the adaptability and stability of the model.
 As shown in Figure 1(b), the proposed method achieves significant improvement against other model update strategies.
-Our contributions can be summarized as follows:
-  • $\bullet$
-    We investigate the test-time adaptation within the realm of VLN. Our proposed method verifies TTA as a promising and viable avenue for enhancing VLN performance.
-  • $\bullet$
-    Based on a unified decomposition-accumulation framework for both gradients and parameters, our method ensures swift model adaptability to environmental changes in the short-term fast update phase, while preserves stability throughout the long-term slow update phase.
-  • $\bullet$
-    Our FSTTA elevates the performance of several leading VLN models across four popular benchmarks.
-When applied to the notable DUET model [11], our method yields a performance boost of over $5\%$ on the representative discrete/continuous datasets REVERIE/R2R-CE.
+Our contributions can be summarized as follows: (1) Considering the characteristics of inter-sample online instruction-execution and intra-sample multi-step action-execution, we explore the online VLN task and propose a fast-slow test-time adaptation (FSTTA) method for effective navigation and model updating.
+(2) Based on a unified decomposition-accumulation framework for both gradients and parameters, our method ensures swift model adaptability to environmental changes in the short-term fast update phase, while preserves stability throughout the long-term slow update phase.
+(3) Our FSTTA elevates the performance of several leading VLN models on four popular benchmarks.
+When applied to the notable DUET model (Chen et al., 2022d), it yields a performance boost of over $5\%$ on the representative discrete/continuous benchmarks REVERIE/R2R-CE.
 Furthermore, our method shows superior results compared to other premier TTA techniques.
@@
 Vision-and-Language Navigation (VLN).
-Recently, VLN tasks have received significant attention in the embodied AI domain and a variety of effective approaches have been proposed [3, 86, 20].
 Most existing methods facilitate VLN research by developing powerful techniques for model training, including: (i) Designing advanced network architectures.
-Sequence-to-sequence framework are the most commonly used one for predicting agent actions from a historical observation series.
-Early VLN models utilize LSTMs with various attention mechanisms [3, 16, 46, 24] while recent ones [23, 25, 9, 41, 11, 27, 1] resort to the more popular transformer-based methods for multi-modal pre-training.
-Other architectures are also explored such as graph neural networks [86] and parameter-efficient adapter [55].
-(ii) Adopting various training paradigms such as reinforcement and imitation learning [49, 63, 74, 17].
-Moreover, to estimate the completeness of instruction following and decide when to conduct backtracking, progress monitoring [45, 85] and back-tracking [30, 46] are also employed to promote training process.
+Early VLN models utilize LSTMs with various attention mechanisms (Anderson et al., 2018; Fried et al., 2018; Hong et al., 2020) while recent ones (Hao et al., 2020; Hong et al., 2021; Chen et al., 2021; Lin et al., 2022; Chen et al., 2022d; Huo et al., 2023) resort to the more popular transformer-based methods for multi-modal pre-training.
+Other architectures are also explored such as graph neural networks (Zhu et al., 2021) and parameter-efficient adapter (Qiao et al., 2023b).
+(ii) Adopting various training paradigms such as reinforcement and imitation learning (Nguyen et al., 2019; Tan et al., 2019; Wang et al., 2019; Gao et al., 2023a).
+Moreover, to estimate the completeness of instruction following and decide when to conduct backtracking, progress monitoring (Ma et al., 2019a; Zhu et al., 2020) and back-tracking (Ke et al., 2019; Ma et al., 2019b) are also employed to promote training process.
 (iii) Performing data augmentation for training a stronger model.
-In recent years, more and more large-scale benchmarks are established via collecting human annotations [33, 86, 57] or creating new environments [52, 10].
-Other approaches explore techniques such as mixup and synthesis [43, 29], style transfer [37], or future-view image semantics [36] for data augmentation.
+In recent years, more and more large-scale benchmarks are established via collecting human annotations (Ku et al., 2020; Zhu et al., 2021; Ramrakhya et al., 2022) or creating new environments (Qi et al., 2020; Chen et al., 2022c).
+Other approaches explore techniques such as mixup and synthesis (Liu et al., 2021; Kamath et al., 2023), style transfer (Li et al., 2022), or future-view image semantics (Li & Bansal, 2023) for data augmentation.
 (iv) Leveraging additional information for boosting model capacity.
-Since the goal of VLN is to navigate in photo-realistic environments, there are many kinds of information in the world that can be used such as knowledge [38], 3D scene geometry [44, 78], and landmarks [72, 12].
-Though the above approaches have attempted to employ fruitful strategies for training an effective VLN model, they still struggle to adequately address the domain discrepancy between training and testing data.
-Some studies [55, 54] have tried to use dynamic networks or auxiliary models to cope with varying test data, however, these approaches neither directly minimize the domain gap between the model and test data nor offer dynamic model update strategies.
-Test-time Adaptation (TTA).
-TTA, which allows models to adapt the test data in an online and unsupervised manner, has gained a lot of interest with a host of various approaches proposed in the literature [39, 40, 61, 73, 34].
-Existing TTA methods generally rely on batch normalization calibration [48, 84, 19], entropy minimization [68, 51, 50, 64], auxiliary self-supervised task or data regularization [62, 28, 5, 83, 66] to acquire useful information for reducing the domain gap between training and testing data.
-To stabilize adaptation in continuously changing data distribution, recently, continual test-time adaptation [71, 50, 60, 6, 13, 82], as a more practical setting, has been tentatively explored for addressing the cumulative errors and catastrophic forgetting issues.
-Until now, test-time adaptation has been preliminarily explored in some sequential data analysis fields such as action recognition [42] and video classification [80].
-However, TTA on VLN tasks is yet to be explored.
+Since the goal of VLN is to navigate in photo-realistic environments, there are many kinds of information in the world that can be used such as knowledge (Li et al., 2023), 3D scene geometry (Liu et al., 2023; Wang et al., 2023f), and landmarks (Wang et al., 2022b; Cui et al., 2023).
+Although the above methods have made significant progress in training effective models, they overlook the utilization of test data during the online VLN process.
+In real-world applications, an agent is required to continuously execute user instructions at different times.
+The ability of an agent to accumulate experience during this process would greatly enhance its practical value.
+Note that (Lu et al., 2022) first explored test-time adaptation for VLN.
+However, this approach does not perform model update in an online manner and overlook the balance between adaptability-stability.
+Online Test-time Adaptation (TTA).
+TTA allows models to adapt the test data in an online and unsupervised manner (Liang et al., 2023; Lim et al., 2023; Lee et al., 2023).
+Existing TTA methods generally rely on batch normalization calibration (Mirza et al., 2022; Zhao et al., 2023; Gong et al., 2022), entropy minimization (Wang et al., 2021; Niu et al., 2023; Tang et al., 2023), auxiliary self-supervised task or data regularization (Sun et al., 2020; Boudiaf et al., 2022; Zhang et al., 2022) to acquire useful information for reducing the domain gap between training and testing data.
+To stabilize adaptation in continuously changing data distribution, recently, continual test-time adaptation (Wang et al., 2022a; Niu et al., 2022; Song et al., 2023; Döbler et al., 2023; Yuan et al., 2023; Liu et al., 2024), as a more practical setting, has been tentatively explored for addressing the cumulative errors and catastrophic forgetting issues.
+Until now, test-time adaptation has been preliminarily explored in some sequential data analysis fields such as action recognition (Lin et al., 2023) and video classification (Yi et al., 2023).
+However, they overlook the joint inter- and intra-sample structure in TTA of sequential data.
 Gradient-based Methods.
-Gradients are typically central to modern SGD-based deep learning algorithms.
-To date, gradient analysis research has predominantly focused on domain generalization (DG) [47, 35, 70, 56, 76, 65], due to the negative impact of conflicting gradients from multiple domains on model optimization.
-Pioneering works [15, 81, 47] perform gradient surgery at the backpropagation phase via various strategies such as normal plane projection [81] and consensus learning [47].
-Other approaches resort to gradient agreement regularization for refining the optimization direction by leveraging sharpness [70] or similarity [58, 56] measurements.
-Different from the above models that only consider a single-phase gradient surgery in DG, we jointly analyze the gradient-parameter states for a two-phase (fast-slow) TTA in the VLN task.
+Gradients are central to modern SGD-based deep learning algorithms.
+To date, gradient analysis research has predominantly focused on domain generalization (DG) (Mansilla et al., 2021; Lew et al., 2023; Wang et al., 2023b; Rame et al., 2022; Wang et al., 2023e; Tian et al., 2023), due to the negative impact of conflicting gradients from multiple domains on model optimization.
+Pioneering works (Du et al., 2018; Yu et al., 2020; Mansilla et al., 2021) perform gradient surgery at the backpropagation phase via various strategies such as normal plane projection (Yu et al., 2020) and consensus learning (Mansilla et al., 2021).
+Other approaches resort to gradient agreement regularization for refining the optimization direction by leveraging sharpness (Wang et al., 2023b) or similarity (Shi et al., 2021; Rame et al., 2022) measurements.
+Different from the above strategies that only consider a single-phase gradient surgery in DG, we jointly analyze the gradient-parameter states for a two-phase (fast-slow) TTA in the VLN task.
 3 Our Approach
-Figure 2: Overall framework of the proposed Fast-Slow Test-Time Adaptation (FSTTA) for VLN tasks.
+Figure 2: Overall framework of the proposed Fast-Slow Test-Time Adaptation (FSTTA) for online VLN.
 In the fast update phase, taking ‘Sample i’ as an example, the model periodically analyzes the gradients ($\{\bm{g}\}$) generated during the recent multi-step navigation and performs a gradient decomposition-accumulation analysis to pinpoint a concordant direction for model update.
@@
 ‘$\text{F}_{1}$’ indicates the first fast update within a test sample.
-3.1 Preliminaries and Framework Overview
 Problem Setup and VLN Base Model.
@@
 During the navigation process, an undirected exploration graph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is progressively constructed, where $\mathcal{V}_{t}$ denotes navigable nodes, $\mathcal{E}_{t}$ indicates the connectivity edges, $t$ is the current timestep.
+Note that a ‘STOP’ node is added to this graph to indicate a stop action, and we connect it with all other nodes.
 At this moment, the agent receives a panoramic view that contains 36 single images.
-The panorama is represented by the image features $\mathcal{R}_{t}$ and their object features $\mathcal{O}_{t}$, where these features can be extracted by pre-trained vision transformers (ViT) [14, 11, 38].
+The panorama is represented by the image features $\mathcal{R}_{t}$ and their object features $\mathcal{O}_{t}$, where these features can be extracted by pre-trained vision transformers (ViT) (Dosovitskiy et al., 2020; Chen et al., 2022d; Li et al., 2023).
 To accomplish the instruction, the agent needs to predict probabilities for the currently navigable nodes and select the most possible one as the next movement action.
@@
 $\bm{s}_{t}=\phi\left(\bm{I},\mathcal{R}_{t},\mathcal{O}_{t},\mathcal{H}_{t};\bm{\Theta}\right),~{}~{}\bm{s}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|}$ (1)
-where $\mathcal{H}_{t}$ indicates the history information that encodes the observed visual features and performed actions [11, 55].
-$\phi(\cdot)$ is the VLN base model such as the dual-scale graph transformer [11, 10], $\bm{\Theta}$ is the learnable model parameters.
+where $\mathcal{H}_{t}$ indicates the history information that encodes the observed visual features and performed actions (Chen et al., 2022d; Qiao et al., 2023b).
+$\phi(\cdot)$ is the VLN base model such as the dual-scale graph transformer (Chen et al., 2022d, c), $\bm{\Theta}$ is the learnable model parameters.
 Framework Overview.
-In this paper, we devote to adjusting the VLN base model during testing process within an unsupervised manner.
+In this paper, we devote to adjusting the VLN base model during testing process within an online and unsupervised manner.
 Our FSTTA framework is illustrated in  Figure 2.
-For each sample, at timestep $t$, we employ the commonly adopted entropy minimization objective [68, 51] for test-time adaption, which aims to reduce the entropy of the probabilities over the current navigable nodes:
+For each sample, at timestep $t$, we employ the commonly adopted entropy minimization objective (Wang et al., 2021; Niu et al., 2023) for test-time adaption, which aims to reduce the entropy of the probabilities over the current navigable nodes:
 $\mathcal{L}({\bm{s}}_{t};\bm{\Theta})=-\sum\nolimits_{i}{\bm{s}}_{t,i}\log({\bm{s}}_{t,i}).$ (2)
@@
 As a result, we only consider a small portion of the model parameters for gradient calculation.
-Since affine parameters in normalization layers capturing data distribution information, numerous TTA methods opt to update these parameters for adaption [68, 51, 39].
-In this paper, we employ the model’s final a few layer-norm operations for TTA and maintain other parameters frozen.
-For brevity, we still use the symbol $\bm{\Theta}$ to represent these parameters to be updated, $\bm{\Theta}\in\mathbb{R}^{D}$.
+Since affine parameters in normalization layers capturing data distribution information, numerous TTA methods opt to update these parameters for adaption (Wang et al., 2021; Niu et al., 2023; Liang et al., 2023).
+In this paper, we employ the model’s a few final layer-norm operations for TTA and maintain other parameters frozen.
+For brevity, we still use the symbol $\bm{\Theta}$ to represent these parameters to be updated in the following sections, $\bm{\Theta}\in\mathbb{R}^{D}$.
 Targeting at fully leveraging the gradient and parameter information, under a unified decomposition-accumulation analysis framework, we propose an effective two-phase adaptation for fast and slow model updates.
 3.1 Fast Update via Gradient Analysis
-At timestep $t$ in the navigation process, the agent is required to select an action (navigable node) by using the predicted score ${\bm{s}}_{t}$.
-With this score, we can calculate the TTA loss (Eq.
+At timestep $t$ in one navigation process, the agent is required to select an action (navigable node) by using the predicted score ${\bm{s}}_{t}$.
+With ${\bm{s}}_{t}$, we can calculate the TTA loss (Eq.
 (2)) and then derive the gradient of the model parameters $\bm{\Theta}$ as: $\bm{g}_{t}=\nabla\mathcal{L}({\bm{s}}_{t};\bm{\Theta})$, $\bm{g}_{t}\in\mathbb{R}^{D}$.
-Traditional TTA methods conduct adaptation independently at each time step, which can exacerbate the issue of cumulative errors [50, 60], particularly in the VLN process that requires frequent action execution.
+Traditional TTA methods conduct adaptation independently at each time step, which can exacerbate the issue of cumulative errors (Niu et al., 2022; Song et al., 2023), particularly in the VLN process that requires frequent action execution.
 Therefore, we propose to conduct a gradient decomposition-accumulation analysis, wherein we periodically analyze the gradients generated during the recent multi-step navigation process and identify a concordant direction for an iteration of model update.
@@
 For the $j$-th update, gradients from previous $M$ steps are collected as $\bm{G}_{j}=\{\widetilde{\bm{g}}_{j,m}\}_{m=1}^{M}$, where $\bm{G}_{j}\in\mathbb{R}^{M\times D}$, $\widetilde{\bm{g}}_{j,m}$ indicates the $t$-th gradient $\bm{g}_{t}$ when $t=M(j-1)+m$.
-Note that these gradients determines the learning direction of our VLN model, and a simple strategy to compute this direction is to take their average $\bar{\bm{g}}_{j}=1/M\sum_{m}\widetilde{\bm{g}}_{j,m}$; however, this inevitably introduce step-specific noise.
+Note that these gradients determine the learning direction of our VLN model, and a simple strategy to compute this direction is to take their average $\bar{\bm{g}}_{j}=1/M\sum_{m}\widetilde{\bm{g}}_{j,m}$; however, this inevitably introduce step-specific noise.
 To avoid the issue, we aim to find a concordant direction among these gradients.
@@
 where $\hat{\bm{G}}_{j}$ is the centered gradient matrix by removing the mean from $\bm{G}_{j}$.
-The $m$-th row vector in $\hat{\bm{G}}_{j}$ reflects the deviation between $\widetilde{\bm{g}}_{j,m}$ and the average gradient $\bar{\bm{g}}_{j}$.
+The $m$-th row in $\hat{\bm{G}}_{j}$ reflects the deviation between $\widetilde{\bm{g}}_{j,m}$ and the average gradient $\bar{\bm{g}}_{j}$.
 $\lambda_{j,d}$, $\bm{u}_{j,d}$ denote the $d$-th largest eigenvalue and the corresponding eigenvector.
-Motivated by the principle component analysis [59], it is obvious that a larger $\lambda_{j,d}$ corresponds to a higher variance of the gradient projection length $\bm{G}_{j}\bm{u}_{j,d}$ and vice versa.
+Motivated by the principle component analysis (Shlens, 2014), it is obvious that a larger $\lambda_{j,d}$ corresponds to a higher variance of the gradient projection length $\bm{G}_{j}\bm{u}_{j,d}$ and vice versa.
 Hence, we can derive a concordant gradient by adaptively aggregating the gradients’ components on all the axes by considering different eigenvalue (importance):
@@
 With $\nabla^{(fast)}$, we can perform fast model update by setting a learning rate $\gamma^{(fast)}$.
-Although traditional methods employ a fixed learning rate during optimization, such a setting might hinder model convergence, i.e., small learning rates slow down convergence while aggressive learning rates prohibit convergence [4].
+Although traditional methods employ a fixed learning rate, such a setting might hinder model convergence, i.e., small learning rates slow down convergence while aggressive learning rates prohibit convergence (Barzilai & Borwein, 1988).
 Since fast updates are frequently invoked during navigation, relying on a fixed learning rate is sub-optimal.
@@
 Different from varying the learning rate through optimizer or scheduler, we argue for a scaling method that leverages gradient agreement information in historical steps to dynamically adjust the speed of model update.
-Current gradient alignment strategies typically impose direct constraints on the gradients [58, 56], which are not suitable for our framework as they undermine the gradient decomposition-accumulation process.
-Given that the second-order information (variance) has been demonstrated to be more effective than the first-order information (mean) in gradient agreement learning [56], we directly utilize the trace of the gradient covariance matrix, $\text{Tr}\left(1/(M-1){\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right)$, for scaling.
+Current gradient alignment strategies typically impose direct constraints on the gradients (Shi et al., 2021; Rame et al., 2022), which are not suitable for our framework as they undermine the gradient decomposition-accumulation process.
+Given that the second-order information (variance) has been demonstrated to be more effective than the first-order information (mean) in gradient agreement learning (Rame et al., 2022), we directly utilize the trace of the gradient covariance matrix, $\text{Tr}\left(1/(M-1){\hat{\bm{G}}}_{j}^{\mathsf{T}}\hat{\bm{G}}_{j}\right)$, for scaling.
 Note that the trace is equal to the sum of eigenvalues $\sigma_{j}=\sum_{d}\lambda_{j,d}$.
@@
 The historical variance $\bar{\sigma}$ is updated as $\bar{\sigma}\leftarrow\rho\bar{\sigma}+(1-\rho)\sigma_{j}$ and maintained for all samples throughout the test stage, $\rho$ is the update momentum.
-Model Update. With the above gradient and learning rate, we can perform the $j$-th fast model update:
+Model Update. With the above gradient and learning rate, we can perform a single iteration as the $j$-th fast update:
 $\bm{\Theta}_{j}=\bm{\Theta}_{j-1}-\gamma_{j}^{(fast)}\cdot\nabla_{j}^{(fast)},$ (7)
@@
 In the fast update phase, although we obtain concordant optimization directions, the frequent parameter updates may still dramatically change the VLN model.
-To maintain the stability of the VLN model during long-term usage, we revert the model to its historical states recorded in the fast update phase, and conduct a decomposition-accumulation analysis on the parameter variation trajectory for direct parameter modulation.
+To maintain the stability of the VLN model during online long-term usage, we periodically revert the model to its historical states, and conduct a decomposition-accumulation analysis on the parameter variation trajectory for direct parameter modulation.
 The slow update phase shares the core formulation with the fast phase, but shifts the focus from gradients to the model parameters themselves.
@@
 Note that in the slow update phase, we additionally incorporate $\bm{\Theta}^{(l-1)}$ from the previous update for analysis since it serves as a starting reference point for direct parameter modulation.
-Table 1: Comparison on REVERIE. Results better than the base model are highlighted in bold.
-Table 2: Results on R2R dataset.
-Methods REVERIE Val Unseen REVERIE Test Unseen
-TL $\downarrow$ OSR SR SPL RGS RGSPL TL $\downarrow$ OSR SR SPL RGS RGSPL
-Human - - - - - - 21.18 86.83 81.51 53.66 77.84 51.44
-Seq2Seq [3] [CVPR18] 11.07 8.07 4.20 2.84 2.16 1.63 10.89 6.88 3.99 3.09 2.00 1.58
-RCM [74] [CVPR19] 11.98 14.23 9.29 6.97 4.89 3.89 10.60 11.68 7.84 6.67 3.67 3.14
-SMNA [45] [ICLR19] 9.07 11.28 8.15 6.44 4.54 3.61 9.23 8.39 5.80 4.53 3.10 2.39
-FAST [52] [CVPR20] 45.28 28.20 14.40 7.19 7.84 4.67 39.05 30.63 19.88 11.61 11.28 6.08
-Airbert [21] [ICCV21] 18.71 34.51 27.89 21.88 18.23 14.18 17.91 34.20 30.28 23.61 16.83 13.28
-HAMT [9] [NeurIPS21] 14.08 36.84 32.95 30.20 18.92 17.28 13.62 33.41 30.40 26.67 14.88 13.08
-HOP [53] [CVPR22] 16.46 36.24 31.78 26.11 18.85 15.73 16.38 33.06 30.17 24.34 17.69 14.34
-LANA [75] [CVPR23] 23.18 52.97 48.31 33.86 32.86 22.77 18.83 57.20 51.72 36.45 32.95 22.85
-BEVBert [1] [ICCV23] - 56.40 51.78 36.37 34.71 24.44 - 57.26 52.81 36.41 32.06 22.09
-BSG [44] [ICCV23] 24.71 58.05 52.12 35.59 35.36 24.24 22.90 62.83 56.45 38.70 33.15 22.34
-GridMM [78] [ICCV23] 23.20 57.48 51.37 36.47 34.57 24.56 19.97 59.55 55.13 36.60 34.87 23.45
-DUET [11] [CVPR22] 22.11 51.07 46.98 33.73 32.15 23.03 21.30 56.91 52.51 36.06 31.88 22.06
-DUET-FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 21.52 58.44 53.40 36.43 32.99 22.40
-HM3D [10] [ECCV22] 22.13 62.11 55.89 40.85 36.58 26.76 20.87 59.81 53.13 38.24 32.69 22.68
-HM3D-FSTTA 22.37 63.74 57.02 41.41 36.97 26.55 21.90 63.68 56.44 39.58 34.05 23.04
-Methods R2R Val Unseen
-TL $\downarrow$ NE $\downarrow$ SR SPL
-Seq2Seq [3] 8.39 7.81 22 -
-RCM [74] 11.46 6.09 43 -
-SMNA [45] - 5.52 45 32
-EnvDrop [63] 10.70 5.22 52 48
-AirBert [21] 11.78 4.10 62 56
-HAMT [9] 11.46 3.65 66 61
-GBE [86] - 5.20 54 43
-SEvol [7] 12.26 3.99 62 57
-HOP [53] 12.27 3.80 64 57
-BEVBert [1] 14.55 2.81 75 64
-LANA [75] 12.00 - 68 62
-BSG [44] 14.90 2.89 74 62
-GridMM [78] 13.27 2.83 75 64
-DUET [11] 13.94 3.31 72 60
-DUET-FSTTA 14.64 3.03 75 62
-HM3D [10] 14.29 2.83 74 62
-HM3D-FSTTA 14.86 2.71 75 63
-Table 2: Results on R2R dataset.
+Table 1: Experimental results for different TTA strategies on REVERIE dataset.
+Methods REVERIE Val Seen REVERIE Val Unseen REVERIE Test Unseen Time(ms)*
+OSR SR SPL RGSPL OSR SR SPL RGSPL OSR SR SPL RGSPL
+DUET (Chen et al., 2022d) 73.86 71.75 63.94 51.14 51.07 46.98 33.73 23.03 56.91 52.51 36.06 22.06 104.84
+   + EATA (Niu et al., 2022) 74.05 72.17 63.41 49.94 52.09 47.40 33.46 22.65 57.21 52.91 35.19 21.65 133.12
+   + CoTTA (Wang et al., 2022a) 73.79 71.05 62.91 49.36 52.46 47.56 31.43 21.83 56.14 52.52 34.66 21.06 3.89$\times$ 103
+   + NOTE (Gong et al., 2022) 74.76 72.43 64.28 51.97 52.85 48.28 33.98 22.98 57.66 53.41 36.18 22.09 137.89
+   + SAR (Niu et al., 2023) 74.84 71.75 64.43 51.70 53.26 48.00 33.92 23.09 57.11 53.04 36.07 22.27 145.53
+   + ViDA (Liu et al., 2024) 73.99 72.49 63.49 50.89 52.53 48.14 32.45 21.92 56.78 52.74 35.10 21.77 3.97$\times$ 103
+   + Tent (Wang et al., 2021) 70.07 70.73 61.67 49.31 49.43 46.87 31.90 20.15 54.58 50.56 33.37 20.32 126.91
+   + Tent-INT-2 71.82 70.36 61.53 49.82 51.22 48.46 33.67 21.30 54.25 50.36 33.89 21.09 124.02
+   + Tent-INT-3 74.54 72.58 64.44 51.05 52.28 48.60 34.65 23.12 56.66 52.92 35.84 21.89 119.34
+   + Tent-INT-4 73.84 72.51 64.03 50.97 51.40 48.91 35.06 22.99 56.74 53.14 36.29 22.14 117.26
+   + Tent-Stable 73.72 71.89 64.06 50.41 51.43 47.55 33.99 23.32 57.12 52.61 36.17 22.16 129.22
+   + FSTTA 75.59 75.48 65.84 52.23 56.26 54.15 36.41 23.56 58.44 53.40 36.43 22.40 135.61
+    Note that the last column displays the average execution time of the agent for a single instruction, calculated on the Validation Unseen set of REVERIE.
+Table 2: Ablation study on REVERIE dataset.
+Module REVERIE Val Unseen
+Fast DLR Slow TL $\downarrow$ OSR SR SPL RGS RGSPL
+- - - 22.11 51.07 46.98 33.73 32.15 23.03
+Tent - - 22.52 52.28 48.60 34.65 32.66 23.12
+- - 22.65 53.50 49.74 34.91 33.70 23.36
+- 22.43 54.01 49.82 35.34 34.32 23.29
+22.14 56.26 54.15 36.41 34.27 23.56
+Table 3: Results on Validation Seen set of REVERIE.
+FSTTA REVERIE Val Seen
+Unseen Seen TL $\downarrow$ OSR SR SPL RGS RGSPL
+ $\times$           $\times$ 13.86 73.86 71.15 63.94 57.41 51.14
+ $\times$  $\longrightarrow$  ✓ 15.13 75.59 75.48 65.84 58.62 52.23
+ ✓  $\longrightarrow$  $\times$ 13.40 73.16 71.78 64.18 57.05 51.18
+ ✓  $\longrightarrow$  ✓ 15.11 75.58 74.12 65.53 59.20 52.18
 Similar to the fast update phase, the centered parameter matrix $\hat{\bm{M}}_{l}$ can be constructed, where the $n$-th row vector in it reflects the deviation between $\widetilde{\bm{\Theta}}_{l,n}$ and the averaged historical parameter $\bar{\bm{\Theta}}_{l}=1/(N+1)\sum_{n}\widetilde{\bm{\Theta}}_{l,n}$.
@@
 $\bm{Z}_{l}=\{{\bm{z}_{l,d}}^{\mathsf{T}}\}_{d=1}^{D}$ depicts the local coordinate system where each axis depicts the direction of parameter variation.
-Intuitively, the principal axes (with larger eigenvalues) delineate the primary directions of historical parameter variation, while minor axes (with smaller eigenvalues) often encompass noise [76].
-To find a more reliable optimization path to traverse the trajectory of primary parameter changes, we pay more attention on the axis with the larger variance.
+Intuitively, the principal axes (with larger eigenvalues) delineate the primary directions of historical parameter variation, while minor axes (with smaller eigenvalues) often encompass noise (Wang et al., 2023e).
+To find a reliable optimization path to traverse the trajectory of primary parameter changes, we pay attention on the axis with the large variance.
 Since there is no silver bullet to learning an optimization direction with only parameters, a reference direction can significantly aid in guiding the model towards a local optimal.
@@
 Besides, the norm of the reference direction is utilized to automatically tuning the magnitude of the analyzed gradient.
-In contrast to $\Phi_{d}(\cdot)$ in gradient analysis, $\Psi_{d}(\cdot)$ highlights those axes with high variation due to the different characteristics of gradients and parameters in model optimization.
+In contrast to $\Phi_{d}(\cdot)$ in the fast update phase, $\Psi_{d}(\cdot)$ highlights those axes with high variation due to the different characteristics of gradients and parameters in model optimization.
 Model Update. With $\nabla_{l}^{(slow)}$, we can perform the $l$-th slow model update as follows:
@@
 where $\gamma^{(slow)}$ is learning rate.
-Since the slow update phase is designed for stable model learning and is not frequently invoked, we employ a fixed learning rate here instead of conducting dynamic learning rate scaling, as done in the fast phase.
-The updated parameter $\bm{\Theta}^{(l)}$ will be utilized for the subsequent test samples in conjunction with new fast update phases applied to them.
+Since the slow update phase is designed for stable model learning and is not frequently invoked, we employ a fixed learning rate here instead of conducting the dynamic learning rate scaling as done in the fast phase.
+The updated parameter $\bm{\Theta}^{(l)}$ will be utilized for the subsequently coming test samples in conjunction with new fast update phases applied to them.
 4 Experimental Results
-We evaluate FSTTA on four benchmarks: REVERIE [52], R2R [3], SOON [86], and R2R-CE [32] datasets. Experiments and ablation studies show our effectiveness.
-4.1 Experimental Setup
+Table 4: Results on Validation Unseen & Seen sets of REVERIE.
+Methods REVERIE Val Unseen & Seen
+TL $\downarrow$ OSR SR SPL RGS RGSPL
+DUET 19.18 61.53 57.49 45.66 41.56 34.38
+   + Tent 20.23 57.33 54.86 41.90 38.09 32.46
+   + EATA 20.29 62.77 57.31 44.59 41.54 34.16
+   + SAR 20.52 63.59 57.80 44.72 42.45 34.88
+   + FSTTA 20.48 63.36 60.23 47.96 43.58 35.65
+Table 5: Experimental results on REVERIE dataset.
+Methods Val Seen Val Unseen Test Unseen
+OSR SR SPL RGSPL OSR SR SPL RGSPL OSR SR SPL RGSPL
+Seq2Seq (Anderson et al., 2018) 35.70 29.59 24.01 14.96 8.07 4.20 2.84 2.16 6.88 3.99 3.09 1.58
+RCM (Wang et al., 2019) 29.44 23.33 21.82 15.36 14.23 9.29 6.97 3.89 11.68 7.84 6.67 3.14
+FAST (Qi et al., 2020) 55.17 50.53 45.50 29.66 28.20 14.40 7.19 4.67 30.63 19.88 11.61 6.08
+RecBERT (Hong et al., 2021) 53.90 51.79 47.96 35.61 35.02 30.67 24.90 15.27 32.91 29.61 23.99 13.51
+Airbert (Guhur et al., 2021) 48.98 47.01 42.34 30.01 34.51 27.89 21.88 14.18 34.20 30.28 23.61 13.28
+HAMT (Chen et al., 2021) 47.65 43.29 40.19 25.18 36.84 32.95 30.20 17.28 33.41 30.40 26.67 13.08
+HOP (Qiao et al., 2022) 53.76 54.88 47.19 33.85 36.24 31.78 26.11 15.73 33.06 30.17 24.34 14.34
+BEVBert (An et al., 2022) 76.18 73.72 65.32 51.73 56.40 51.78 36.37 24.44 57.26 52.81 36.41 22.09
+LANA (Wang et al., 2023d) 74.28 71.94 62.77 50.34 52.97 48.31 33.86 22.77 57.20 51.72 36.45 22.85
+BSG (Liu et al., 2023) 78.36 76.18 66.69 54.02 58.05 52.12 35.59 24.24 62.83 56.45 38.70 22.34
+GridMM (Wang et al., 2023f) - - - - 57.48 51.37 36.47 24.56 59.55 55.13 36.60 23.45
+DUET (Chen et al., 2022d) 73.86 71.75 63.94 51.14 51.07 46.98 33.73 23.03 56.91 52.51 36.06 22.06
+DUET-FSTTA 75.59 75.48 65.84 52.23 56.26 54.15 36.41 23.56 58.44 53.40 36.43 22.40
+HM3D (Chen et al., 2022c) 66.76 65.00 55.70 41.66 62.11 55.89 40.85 26.76 59.81 53.13 38.24 22.68
+HM3D-FSTTA 69.41 67.79 58.28 41.60 63.74 57.02 41.41 26.55 63.68 56.44 39.58 23.04
 Datasets.
-Four datasets are adopted for conducting our experiments.
-Among them, REVERIE [52] contains 10,567 panoramic images and 21,702 high-level instructions, focusing on grounding remote target object within 90 buildings.
-R2R [3] provides step-by-step instructions for navigation in photo-realistic environments, which includes 10,800 panoramic views and 7,189 trajectories.
-SOON [86] also requires the agent to find the target object with a more detailed description of the goal.
+We use the popular and standard VLN benchmark REVERIE (Qi et al., 2020) to investigate test-time adaptation within the realm of online VLN.
+REVERIE contains 10,567 panoramic images and 21,702 high-level instructions, focusing on grounding remote target object within 90 buildings.
+In addition, we also adopt other three benchmarks for evaluating the effectiveness of our proposed FSTTA.
+Among them, R2R (Anderson et al., 2018) provides step-by-step instructions for navigation in photo-realistic environments, which includes 10,800 panoramic views and 7,189 trajectories.
+SOON (Zhu et al., 2021) also requires the agent to find the target object with a more detailed description of the goal.
 It has 3,848 sets of instruction and more than 30K long distance trajectories.
-R2R-CE [32] is a variant of R2R in continuous environments, where an agent is able to move freely and engage with obstacles.
+Note that the performance comparisons are based on their challenge report, specifically within the provided val unseen and test unseen splits.
+R2R-CE (Krantz et al., 2020) is a variant of R2R in continuous environments, where an agent is able to move freely and engage with obstacles.
 The dataset consists of 16,000 instruction-trajectory pairs, with non-transferrable paths excluded.
 Evaluation Metrics.
-We follow previous approaches [52, 11, 10, 37, 78] and employ the most commonly used metrics for evaluating VLN agents as follows: TL (Trajectory Length), NE (Navigation Error), SR (Success Rate), SPL (Success weighted by Path Length), OSR (Oracle Success Rate), RGS (Remote Grounding Success rate), and RGSPL (RGS weighted by Path Length).
+We follow previous approaches (Qi et al., 2020; Chen et al., 2022d, c; Li et al., 2022; Wang et al., 2023f) and employ the most commonly used metrics for evaluating VLN agents, i.e., TL (Trajectory Length): the agent’s average path length in meters; NE (Navigation Error): average distance in meters between the agent’s final location and the target one; SR (Success Rate): the proportion of successfully executed instructions with the NE less than 3 meters; SPL (Success weighted by Path Length): SR penalized by Path Length, which is calculated as $\frac{1}{E}\sum_{i=1}^{E}S_{i}\frac{l_{i}}{max(p_{i},l_{i})}$, where E is the number of tasks, $S_{i}$ denotes the success as a binary value, $l_{i}$ and $p_{i}$ denote the shortest path and actual path length for the $i^{th}$ task; OSR (Oracle Success Rate): SR given the oracle stop policy; RGS (Remote Grounding Success rate): proportion of successfully executed instructions where the output bounding box has an IoU (intersection over union) $\geq$ 0.5 with the ground truth; and RGSPL (RGS weighted by Path Length): RGS penalized by Path Length.
+Among them, SR and SPL are the most common metrics for evaluation.
 Implementation Details.
-To better conform to practical scenarios, for all datasets, we set the batch size to 1 during evaluation.
-Each sample (or each action step) is forward propagated only once during the testing process.
-We adopt DUET [11] and HM3D [10] as the base models.
-Since HM3D does not provide training code for R2R-CE dataset, we adopt another state-of-the-art method, BEVBert [1], for TTA.
-Note that for the base models, in Section 4.2, we report the results obtained from running their official codes.
-For VLN models equipped with TTA strategies, we run the corresponding experiments 5 times while shuffling the order of the samples and report the average results.
+To better conform to practical applications, we set batch size to 1 during evaluation, where each sample (and each action step) is forward propagated only once.
+Owing to the lack of an authentic online VLN evaluation setting, we shuffle test samples in each dataset split and sequentially input them into the agent to simulate the online execution and adaptation.
+Specifically, for VLN models equipped with TTA strategies, we run the experiments with shuffled samples 5 times and report the average results.
+We adopt DUET (Chen et al., 2022d) and HM3D (Chen et al., 2022c) as the base models.
+Since HM3D does not provide training code for R2R-CE, we adopt another SOTA methods, WS-MGMap (Chen et al., 2022b) and BEVBert (An et al., 2022), for TTA.
+Note that for the base models, we report the results obtained from running their official codes.
 In our FSTTA, we only utilize the last four LN layers of base models for model updating, all the feature dimensions of these layers are 768.
@@
 All experiments are conducted on a RTX 3090 GPU.
+4.1 Comparison with Different TTA Strategies
+Currently, various TTA methods have been adeptly integrated for the dynamic model updates, marking significant progress.
+Although the exploration of TTA within the VLN field remains relatively untapped, the integration of contemporary advanced TTA methodologies into VLN is feasible.
+Since efficiency is an important evaluation metric for TTA, we provide the average time taken by each method to execute a single instruction for comparison.
+For the compared methods, SAR and TENT are the popular entropy minimization models, whereas NOTE, CoTTA, EATA and ViDA are state-of-the-art continual TTA methods.
+The results in Table 1 demonstrate the capability of our proposed FSTTA to blend model performance with testing efficiency.
+Specifically, on the validation unseen dataset, our method exhibits a discernible enhancement of 6.2% and 2.5% on the SR and SPL metrics compared to the state-of-the-art SAR method, concurrently manifesting a reduction of 7% in testing time.
+From the results, we observe that directly applying existing TTA methods to the online VLN task does not lead to significant performance improvements.
+Furthermore, we investigate different frequencies of updates based on TENT as well as the stable update approach.
+‘INT’ denotes the update interval, which means averaging the gradient information over a certain interval and then performing an iteration of model update; these results are consistent with those in Figure 1(b).
+It can be seen that our method still outperforms these strategies with marginally increased time costs.
+Note that only the optimal values of experimental results are highlighted in bold across all tables.
+Moreover, from Tables 5 to 8, different font colors are employed to indicate whether our method exceeds the performance of the corresponding baseline methods, i.e., using red to denote superior results and blue for inferior ones.
+Table 6: Experimental results on R2R dataset.
+Methods Val Seen Val Unseen
+TL $\downarrow$ NE $\downarrow$ SR SPL TL $\downarrow$ NE $\downarrow$ SR SPL
+Seq2Seq (Anderson et al., 2018) 11.33 6.01 39 - 8.39 7.81 22 -
+RCM (Wang et al., 2019) 10.65 3.53 67 - 11.46 6.09 43 -
+EnvDrop (Tan et al., 2019) 11.00 3.99 62 59 10.70 5.22 52 48
+PREVALENT (Hao et al., 2020) 10.32 3.67 69 65 10.19 4.71 58 53
+RecBERT (Hong et al., 2021) 11.13 2.90 72 68 12.01 3.93 63 57
+HAMT (Chen et al., 2021) 11.15 2.51 76 72 11.46 3.65 66 61
+HOP (Qiao et al., 2022) 11.26 2.72 75 70 12.27 3.80 64 57
+DAVIS (Lu et al., 2022) 12.45 3.16 80 76 12.65 3.16 67 61
+BEVBert (An et al., 2022) 13.56 2.17 81 74 14.55 2.81 75 64
+DUET (Chen et al., 2022d) 12.33 2.28 79 73 13.94 3.31 72 60
+DUET-FSTTA 13.39 2.25 79 73 14.64 3.03 75 62
+HM3D (Chen et al., 2022c) 13.30 2.70 77 71 14.29 2.83 74 62
+HM3D-FSTTA 13.52 2.57 78 71 14.86 2.71 75 63
+Table 7: Experimental results on SOON dataset.
+Methods Val Unseen Test Unseen
+OSR SR SPL RGSPL OSR SR SPL RGSPL
+GBE (Zhu et al., 2021) 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
+GridMM (Wang et al., 2023f) 53.39 37.46 24.81 3.91 48.02 36.27 21.25 4.15
+KERM (Li et al., 2023) 51.62 38.05 23.16 4.04 - - - -
+AZHP (Gao et al., 2023a) 56.19 40.71 26.58 5.53 - - - -
+DUET (Chen et al., 2022d) 50.91 36.28 22.58 3.75 43.00 33.44 21.42 4.17
+DUET-FSTTA 52.57 36.53 23.82 3.75 43.44 35.34 23.23 4.52
+HM3D (Chen et al., 2022c) 53.22 41.00 30.69 4.06 47.26 40.26 28.09 5.15
+HM3D-FSTTA 54.19 42.44 31.03 4.93 48.52 42.02 28.95 5.20
+4.2 Extensive Analysis of FSTTA
+Ablation Studies of the Proposed FSTTA.
+In this work, we propose a FSTTA method for online VLN, which consists of both fast and slow model update phases.
+To validate their effectiveness, we progressively integrate the two phases into the baseline DUET model.
+In addition, we design a baseline variant, which equips DUET with the vanilla TTA objective (TENT (Wang et al., 2021)) and simply utilize the averaged gradient in an interval (with the same $M$) for fast model updates.
+Empirical findings from Table 2 illuminate that the integration of fast and slow phases progressively bolsters the base model by 2.8% and 4.3% on the SR metric.
+Moreover, the dynamic learning rate scaling module (DLR) also contributes to enhancing the model’s performance.
+Will our method experience catastrophic forgetting?
+For an online VLN agent endowed with the TTA capability, it faces the issue of catastrophic forgetting of historical environments and instructions upon continually executing new instructions in new environments.
+To assess whether our method harbors this issue, we re-evaluate our methods on the REVERIE validation seen set.
+Compared with the base model, as shown in Table 3, we find that: (1) Obviously, directly applying FSTTA with the base model on seen data can noticeably enhance performance.
+(2) After performing FSTTA on the unseen set, the obtained model, when tested directly on the seen dataset without TTA, achieves performance comparable to the base model, confirming that our method does not suffer from catastrophic forgetting.
+(3) Applying the updated model from unseen set to the seen set with TTA yields comparable results against the seen-only-TTA version.
+This indicates that our method is effective in environment adaption and experience accumulation.
+Generalization Testing in More Practical Environments.
+In practical applications, agents might encounter both previously seen and unseen scenarios.
+In preceding experiments, we exclusively test on the validation seen and unseen sets separately.
+To verify the generalizability, we combine the seen and unseen sets into a unified set for online VLN.
+Table 4 shows that FSTTA outperforms other TTA methods in effectively managing a variety of testing scenarios.
+Table 8: Experimental results on R2R-CE dataset.
+Methods Val Unseen Test Unseen
+NE $\downarrow$ OSR SR SPL NE $\downarrow$ OSR SR SPL
+Seq2Seq (Krantz et al., 2020) 7.37 40 32 30 7.91 36 28 25
+CM2 (Georgakis et al., 2022) 7.02 42 34 28 7.70 39 31 24
+Sim2Sim (Krantz & Lee, 2022) 6.07 52 43 36 6.17 52 44 37
+CWP-BERT (Hong et al., 2022) 5.74 53 44 39 5.89 51 42 36
+DREAMW (Wang et al., 2023a) 5.53 49 59 44 5.48 49 57 44
+GridMM (Wang et al., 2023f) 5.11 61 49 41 5.64 56 46 39
+ETPNav (An et al., 2023) 4.71 65 57 49 5.12 63 55 48
+WS-MGMap (Chen et al., 2022b) 6.28 48 39 34 7.11 45 35 28
+WS-MGMap-FSTTA 6.16 49 40 35 7.62 46 37 28
+DUET (Chen et al., 2022d) 5.13 55 46 40 5.82 50 42 36
+DUET-FSTTA 5.27 58 48 42 5.84 55 46 38
+BEVBert (An et al., 2022) 4.57 67 59 50 4.70 67 59 50
+BEVBert-FSTTA 4.39 65 60 51 5.45 69 60 50
 4.3 Comparison with State-of-the-art VLN Models
 REVERIE.
-Table 2 presents a comparison of our FSTTA against state-of-the-art methods on the REVERIE dataset.
-Compared with the base models which do not perform test-time adaptation, the proposed method demonstrates favorable performance improvement across most evaluation metrics across the two dataset splits.
-Specifically, on the validation unseen split, our model exhibits notable advantages over DUET, with improvements of 5.3% on OSR, 7.1% on SR, and 2.7% on SPL.
-Furthermore, for the recent state-of-the-art method HM3D, our model displays enhanced generalization capabilities on the test unseen split, achieving remarkable improvements over HM3D, including 3.9%, 3.3%, and 1.3% increases on the three metrics.
-Compared with other state-of-the-arts, our proposed method can achieve superior or comparable performance.
+Table 5 presents a comparison on REVERIE dataset.
+Compared with the base models which do not perform test-time adaptation, the proposed method shows favorable performance improvement across most evaluation metrics across the three dataset splits.
+Specifically, on the validation unseen split, our model exhibits notable advantages over DUET, with improvements of 7.1% on SR, and 2.7% on SPL.
 These results unequivocally affirm the effectiveness of our fast-slow test time adaptation model, showing the promising potential of TTA in the VLN field.
-It is noteworthy that none of the prior methods employed a TTA strategy on this task.
 R2R.
@@
 We speculate that a possible reason is that performing TTA online may increase the likelihood of the agent deviating from its original action execution pattern, leading to more exploration or backtracking.
-This situation is further confirmed in the analysis of various TTA strategies in Table 5.
+This situation is also confirmed in the analysis of various TTA strategies in Table 1.
 SOON.
@@
 On the test unseen split, our approach improves the performance of DUET by substantial gains (e.g., $21.42\%\rightarrow 23.23\%$ for SPL).
-Table 3: Experimental results on SOON dataset.
-Methods Val Unseen Test Unseen
-OSR SR SPL RGSPL OSR SR SPL RGSPL
-GBE [86] 28.54 19.52 13.34 1.16 21.45 12.90 9.23 0.45
-GridMM [78] 53.39 37.46 24.81 3.91 48.02 36.27 21.25 4.15
-DUET [11] 50.91 36.28 22.58 3.75 43.00 33.44 21.42 4.17
-DUET-FSTTA 52.57 36.53 23.82 3.75 43.44 35.34 23.23 4.52
-HM3D [10] 53.22 41.00 30.69 4.06 47.26 40.26 28.09 5.15
-HM3D-FSTTA 54.19 42.44 31.03 4.93 48.52 42.02 28.95 5.20
 R2R-CE.
 FSTTA also generalizes well on the continuous environment, i.e., R2R-CE dataset, as shown in Table 8.
-The results indicate that our approach demonstrates superior or comparable performance against other methods across several metrics.
-Table 4: Experimental results on R2R-CE dataset.
-Methods Val Unseen Test Unseen
-NE $\downarrow$ OSR SR SPL NE $\downarrow$ OSR SR SPL
-Seq2Seq [32] 7.37 40 32 30 7.91 36 28 25
-CWTP [8] 7.90 38 26 23 - - - -
-[…]
-Sim2Sim [31] 6.07 52 43 36 6.17 52 44 37
-CWP-BERT [26] 5.74 53 44 39 5.89 51 42 36
-DREAMW [69] 5.53 49 59 44 5.48 49 57 44
-GridMM [78] 5.11 61 49 41 5.64 56 46 39
-ETPNav [2] 4.71 65 57 49 5.12 63 55 48
-DUET [11] 5.13 55 46 40 5.82 50 42 36
-DUET-FSTTA 5.27 58 48 42 5.84 55 46 38
-BEVBert [1] 4.57 67 59 50 4.70 67 59 50
-BEVBert-FSTTA 4.39 65 60 51 5.45 69 60 50
-Table 5: Experimental results for different TTA strategies.
-Methods REVERIE Val Unseen Time(ms)
-TL $\downarrow$ OSR SR SPL RGS RGSPL
-DUET [11] 22.11 51.07 46.98 33.73 32.15 23.03 104.84
- + EATA [50] 23.41 52.09 47.40 33.46 32.09 22.65 133.12
-[…]
- + NOTE [19] 23.15 52.85 48.28 33.98 32.77 22.98 137.89
- + SAR [51] 23.47 53.26 48.00 33.92 33.49 23.09 145.53
- + Tent [68] 24.05 49.43 46.87 31.90 30.04 20.15 126.91
- + Tent-INT-2 24.24 51.22 48.46 33.67 32.43 21.30 124.02
- + Tent-INT-3 22.52 52.28 48.60 34.65 32.66 23.12 119.34
- + Tent-INT-4 22.59 51.40 48.91 35.06 32.59 22.99 117.26
- + Tent-Stable 22.05 51.43 47.55 33.99 32.34 23.32 129.22
- + FSTTA 22.14 56.26 54.15 36.41 34.27 23.56 135.61
-4.3 Results for Different TTA Strategies
-Currently, various TTA methods have been adeptly integrated for the dynamic model updates within diverse computer vision tasks, marking significant progress.
-Although the exploration of TTA’s application within the VLN field remains relatively untapped, the integration of contemporary advanced TTA methodologies into VLN is feasible.
-Since efficiency is an important evaluation metric for TTA, we provide the average time taken by each method to execute a single instruction for comparison.
-Obviously, equipping with TTA inevitably incurs additional time costs.
-For the compared methods, SAR and TENT are the popular entropy minimization models, whereas NOTE, CoTTA, and EATA are state-of-the-art continual TTA methods.
-The results in Table 5 demonstrate the capability of our proposed FSTTA to blend model performance with testing efficiency.
-Specifically, on the validation unseen dataset of REVERIE, our method exhibits a discernible enhancement of 6.2% and 2.5% on the SR and SPL metrics compared to the state-of-the-art SAR method, concurrently manifesting a reduction of 7% in testing time.
-From the results, directly applying existing TTA methods to the VLN task does not lead to significant performance improvements.
-Furthermore, we investigate different frequencies of updates based on TENT as well as the stable update approach.
-‘INT’ represents the update interval, which means averaging the gradient information over a certain interval and then performing an iteration of model update; these results are consistent with those in Figure 1(b).
-It can be seen that our method still outperforms these strategies with marginally increased time costs.
-Table 6: Ablation study on REVERIE dataset.
-Module REVERIE Val Unseen
-Fast DLR Slow TL $\downarrow$ OSR SR SPL RGS RGSPL
-- - - 22.11 51.07 46.98 33.73 32.15 23.03
-Tent - - 22.52 52.28 48.60 34.65 32.66 23.12
-- - 22.65 53.50 49.74 34.91 33.70 23.36
-- 22.43 54.01 49.82 35.34 34.32 23.29
-22.14 56.26 54.15 36.41 34.27 23.56
-Table 7: Results on validation seen set of REVERIE.
-FSTTA REVERIE Val Seen
-Unseen Seen TL $\downarrow$ OSR SR SPL RGS RGSPL
-- - 13.86 73.86 71.15 63.94 57.41 51.14
-- 15.13 75.59 75.48 65.84 58.62 52.23
-- 13.40 73.16 71.78 64.18 57.05 51.18
-15.11 75.58 74.12 65.53 59.20 52.18
-4.4 Further Remarks
-We perform ablation studies and other in-depth analysis of FSTTA on the validation unseen set of REVERIE [52].
-Ablation Studies of the Proposed FSTTA.
-In this work, we propose a FSTTA method for vision-and-language navigation, which consists of both fast and slow model update phases.
-To validate their effectiveness, we progressively integrate the two phases into the baseline DUET model.
-In addition, we design a baseline variant, which equips DUET with the vanilla TTA objective (TENT [68]) and simply utilize the averaged gradient in an interval (with the same $M$) for fast model updates.
-Empirical findings from Table 6 illuminate that the integration of fast and slow phases progressively bolsters the base model by 2.8% and 4.3% on the SR metric.
-Moreover, the dynamic learning rate scaling module (DLR) also contributes to enhancing the model’s performance.
-Furthermore, our method surpasses the vallina TTA method by a significant margin, showing the consideration of the fast-slow update mechanism is effective.
-Will our method experience catastrophic forgetting?
-For a VLN agent endowed with the TTA capability, it faces the issue of catastrophic forgetting of historical environments and instructions upon continually executing new instructions in new environments.
-To assess whether our method harbors this issue, we re-evaluate our methods on REVERIE validation seen data.
-Compared with the base model, as shown in Table 7, we find that: (1) Directly applying FSTTA with the base model on seen data can noticeably enhance performance.
-(2) After performing FSTTA on the unseen set, the obtained model, when tested directly on the seen dataset without TTA, achieves performance comparable to the base model, confirming that our method does not suffer from catastrophic forgetting.
-(3) Applying the updated model from unseen set to the seen set with TTA yields the best results.
-This indicates that our method is effective in accumulating experience from historical test data.
-Generalization Testing in More Practical Environments.
-In the real-world applications, agents might encounter both previously seen and unseen scenarios.
-In our preceding experiments, we exclusively test on the validation seen and unseen sets separately.
-To verify the generalizability, we combine the seen and unseen sets into a unified set.
-Table 8 shows that FSTTA outperforms other TTA methods in effectively managing a variety of testing scenarios.
+The results indicate that our approach demonstrates superior or comparable performance against other methods.
 Qualitative Analysis.
 Figure 3 provides a visualization of the agent’s instruction execution process, validating that our proposed FSTTA approach can indeed dynamically enhance the VLN performance of the agent during testing.
-Table 8: Results of validation unseen & seen on REVERIE dataset.
-Methods REVERIE Val Unseen & Seen
-TL $\downarrow$ OSR SR SPL RGS RGSPL
-DUET [11] 19.18 61.53 57.49 45.66 41.56 34.38
- + Tent [68] 20.23 57.33 54.86 41.90 38.09 32.46
- + EATA [50] 20.29 62.77 57.31 44.59 41.54 34.16
- + SAR [51] 20.52 63.59 57.80 44.72 42.45 34.88
- + FSTTA 20.48 63.36 60.23 47.96 43.58 35.65
-Figure 3: Visual results of DUET [11] and FSTTA on REVERIE validation unseen set. Our method arrives the correct endpoint.
+Figure 3: Representative visual results on REVERIE validation unseen set.
+Yellow points denote start locations, while the directed lines with red and green points depict the predicted trajectories with target and incorrect endpoints, respectively.
+With FSTTA, the basic agent (DUET) demonstrates enhanced exploration capabilities, effectively moving towards the correct direction, and succeeds based on the object context and scene layouts.
 5 Conclusions
-This paper explores the feasibility of TTA strategies in VLN.
+This paper explores the feasibility of TTA strategies for online VLN.
 We propose a fast-slow test-time adaptation method, which performs decomposition-accumulation analysis for both gradients and parameters, achieving a balance between adaptability and stability.
 The encouraging performance is validated in extensive experiments.
-Several limitations of this paper are noteworthy.
-Firstly, our approach focuses on adapting normalization layers within the trained model.
-While normalization layers are widely employed in deep learning, there are still a few methods that do not utilize these settings.
+Limitations.
+Several limitations are noteworthy.
+Firstly, our approach focuses on adapting normalization layers within the model.
+While these layers are widely employed in deep learning, there are still a few methods that do not utilize them.
 One viable approach to address this issue is to introduce additional normalization layers to the corresponding models and retrain them using the training data.
 In the future, we will also explore how our model can update other types of layers.
-Secondly, the VLN task itself is a cross-modal learning task.
-However, our TTA process does not explicitly consider this information.
-We plan to consider cross-modal TTA in the future.
+Secondly, In this paper, we simulate the online VLN setting simply by sequentially inputting data from the test set.
+In the future, we aim to construct a more realistic agent online learning dataset that aligns with practical application scenarios, to better evaluate TTA methods.
 Thirdly, compared to the base model, the introduction of TTA inevitably incurs additional computational cost, which is a direction for future improvement.